### 20 Mar 2026 v1

# Two-Time-Scale Learning Dynamics: A Population View of Neural Network Training

Abstract
Population-based learning paradigms, including evolutionary strategies, Population-Based Training (PBT), and recent model-merging methods, combine fast within-model optimisation with slower population-level adaptation. Despite their empirical success, a general mathematical description of the resulting collective training dynamics remains incomplete. We introduce a theoretical framework for neural network training based on two-time-scale population dynamics. We model a population of neural networks as an interacting agent system in which network parameters evolve through fast noisy gradient updates of SGD/Langevin type, while hyperparameters evolve through slower selection–mutation dynamics. We prove the large-population limit for the joint distribution of parameters and hyperparameters and, under strong time-scale separation, derive a selection–mutation equation for the hyperparameter density. For each fixed hyperparameter, the fast parameter dynamics relaxes to a Boltzmann–Gibbs measure, inducing an effective fitness for the slow evolution. The averaged dynamics connects population-based learning with bilevel optimisation and classical replicator–mutator models, yields conditions under which the population mean moves toward the fittest hyperparameter, and clarifies the role of noise and diversity in balancing optimisation and exploration. Numerical experiments illustrate both the large-population regime and the reduced two-time-scale dynamics, and indicate that access to the effective fitness, either in closed form or through population-level estimation, can improve population-level updates.
Keywords: Population-based training, hyperparameter optimisation, multi-agent system, propagation of chaos, multi-scale dynamics
1 Introduction
Modern large-scale machine learning increasingly resembles the evolution of model populations rather than the training of a single neural network. In reinforcement learning, hyperparameter optimisation, neural architecture search, and model evolution pipelines, one repeatedly encounters systems in which multiple agents (neural networks, trajectories, or strategies) coexist, interact, and adapt over time. Examples include evolutionary strategies for black-box optimisation (Salimans et al., 2017; Cui et al., 2018), population-based reinforcement learning as in AlphaStar (Vinyals et al., 2019), and recent population-level perspectives on collective learning dynamics (Bouteiller et al., 2026). Further examples arise in distributed AutoML frameworks and in more recent model-merging and evolution methods that improve networks by recombining existing ones (Akiba et al., 2025; Wortsman et al., 2022). These developments suggest that modern AI training pipelines cannot be understood solely in terms of single-trajectory optimisation, but require a population-level perspective based on interaction rules and evolutionary mechanisms.
Within this broader context, hyperparameter optimisation (HPO) has long been recognised as a central component of efficient machine learning pipelines. Classical approaches such as grid search and random search (Bergstra and Bengio, 2012) explore fixed parameter configurations, while Bayesian optimisation methods (Bergstra et al., 2011; Hutter et al., 2011) introduce surrogate models to balance exploration and exploitation. Multi-fidelity strategies, including successive halving and Hyperband (Li et al., 2018), further reduce computational cost by adaptively pruning poor configurations. While effective in many settings, these methods typically operate on independent trials and do not explicitly model interaction or adaptation across agents.
Population-based methods form a distinct paradigm within HPO and meta-optimisation. Among them, Population-Based Training (PBT), originally introduced by DeepMind (Jaderberg et al., 2017; Li et al., 2019a), maintains a population of neural networks that evolve through alternating phases of gradient-based learning and fitness-driven selection. Underperforming agents are periodically replaced by copies of better ones, often with perturbed hyperparameters, enabling the joint optimisation of model parameters and hyperparameters within a single training loop. Variants of this mechanism appear in evolutionary reinforcement learning, quality-diversity optimisation, and large-scale meta-optimisation, and are closely related to evolutionary strategies and population-driven metaheuristics (Dalibard and Jaderberg, 2021; Salimans et al., 2017; Lee et al., 2025; Zito et al., 2025; Hutter et al., 2019). Comprehensive surveys of HPO trends and methodologies can be found in (Franceschi et al., 2025).
Beyond explicitly evolutionary algorithms, similar population-level dynamics arise implicitly in modern large-scale learning systems. In particular, the success of ensemble-based training and selection strategies in systems such as AlphaFold (Jumper et al., 2021) highlights how effective learning can emerge from the interaction and aggregation of multiple trained models, even when the underlying algorithm is not formulated explicitly in evolutionary terms. More recently, classical PBT replacement mechanisms have been extended to act directly on model weights: instead of replace-and-mutate updates, models may be merged through learned recombination rules (Akiba et al., 2025) or by simple averaging of fine-tuned checkpoints (Wortsman et al., 2022). Taken together, these examples reveal a gap in current theory, namely the lack of a mathematical framework able to explain and analyse the collective dynamics generated by interacting populations of models.
From an optimisation perspective, population-based training can also be interpreted as a particular instance of bilevel optimisation (Franceschi et al., 2018; Trillos et al., 2024), where an inner problem performs short gradient-based updates on model parameters, while an outer mechanism selects and adapts hyperparameters or structural traits. This viewpoint naturally induces a separation of time scales and clarifies the algorithmic role of selection and mutation, but still leaves open the question of how to describe such coupled dynamics in a principled continuous-time setting.
Neural Network agents
$(\theta^{i}(t),h^{i}(t)),\,i=1,\dots,N$
$\theta^{i}(t)\leftarrow$ parameters training via SGD/Adam/… $h^{i}(t)\leftarrow$ hyperparameters evolutionary update
Kinetic PDE
Density $f=f(t,\theta,h)$
$\mathcal{L}(\theta,h)$: loss function $\mathcal{F}(\theta,h)$: fitness function $\hskip 42.67912pt\displaystyle\frac{\partial f}{\partial t}=\frac{1}{\varepsilon}\,T_{\mathcal{L}}[f]\;+\;E_{\mathcal{F}}[f]$
Reduced NNs dynamics
$h^{i}(t)\,,\;i=1,\dots,N$
$\theta^{i}(t)\sim\mu^{\infty}(\theta\,|\,h^{i}(t))$ sampling/time averaging $h^{i}(t)\leftarrow$ evolutionary update with effective fitness
Averaged PDE
Ansatz $f(t,\theta,h)=\rho(t,h)\,\mu_{\mathcal{L}}^{\infty}(\theta\,|\,h)$
$\overline{\mathcal{F}}(h)\!=\!\log\int e^{\mathcal{F}(\theta,h)}\mu_{\mathcal{L}}^{\infty}(\theta|h)d\theta$: effective fitness $\frac{\partial\rho}{\partial t}=\overline{E}_{\overline{\mathcal{F}}}[\rho]$
large NN[Section 2]large NNtwo-scale trainingtwo-scale separation ($\varepsilon\to 0$)[Section 3]Refer to captionRefer to caption
Figure 1: Two-scale modeling pipeline. Large-population limits ($N\to\infty$) connect the agent-based neural-network dynamics to PDE descriptions; a subsequent two-scale separation yields reduced hyperparameter dynamics and an averaged PDE.
Despite its conceptual appeal and practical impact, PBT still lacks a rigorous continuous mathematical formulation. The original method is algorithmic, heuristic, and strongly dependent on implementation details such as mutation schedules, population size, and replacement policies. Yet its underlying principles, selection, mutation, adaptation, and interaction, are deeply connected to well-established mathematical theories, including interacting particle systems, kinetic equations, and mean-field limits. Recent works have already begun to explore this direction for related swarm-based optimisation methods (Duong et al., 2020; Borghi and Pareschi, 2025; Borghi et al., 2024; Fornasier et al., 2021; Carrillo et al., 2024; Benfenati et al., 2022; Carrillo et al., 2021; Lu et al., 2024), showing that kinetic and mean-field modelling provides a powerful language for capturing high-dimensional optimisation and evolutionary search. This is different from the mean-field modelling considered in (Mei et al., 2018; Sirignano and Spiliopoulos, 2020), where the training of a single neural network is studied in the limit of infinitely many parameters.
The goal of this manuscript is to provide a rigorous continuous-time theory for neural populations undergoing joint optimisation and evolution. We introduce a continuous-time framework for large-population neural training in which agents undergo gradient learning, selection, and mutation.
From a theoretical point of view, minibatch stochastic gradient descent admits a well-known interpretation as a discretised Langevin diffusion (Mandt et al., 2017; Li et al., 2019b), whose stationary distribution is asymptotically Gibbsian (Chaudhari and Soatto, 2018). This observation provides a natural foundation for two-time-scale modelling approaches, in which fast stochastic gradient dynamics rapidly equilibrate model parameters, while slower population-level mechanisms act on hyperparameters or higher-level traits. At a macroscopic level, such dynamics are closely related to classical selection–mutation models and population transport equations (Ackleh et al., 2016; Alfaro and Veruete, 2019; Bencheikh and Jourdain, 2022; Del Moral and Miclo, 2001; Del Moral, 2001; Perthame, 2007). This separation enables a mathematically tractable description, clarifies the role of diversity and stability, and connects evolutionary heuristics to replicator–mutation dynamics and transport equations.
Our main contributions are:
  • Large-population limit (Theorem 4). We introduce a continuous-time interacting agent model for PBT in which network parameters evolve via Langevin/SGD dynamics and hyperparameters evolve via selection–mutation. We rigorously prove that the multi-agent system converges to a kinetic PDE description as the population size $N\to\infty$ via coupling techniques and a truncated Wasserstein distance.
  • Two-time-scale reduction (Section 3). Under strong time-scale separation ($\varepsilon\to 0$), we derive a closed selection–mutation PDE for the hyperparameter density $\rho(t,h)$, driven by an effective fitness $\overline{\mathcal{F}}(h)$ that averages the fitness against the Boltzmann–Gibbs equilibrium of the fast parameter dynamics. We show this reduction connects PBT to bilevel optimisation (Lemma 8) and to classical replicator–mutator equations from mathematical biology (Section 3.2).
  • Convergence guarantee (Theorem 10). We prove that, under a uniqueness assumption on the effective fitness maximiser, the population mean converges toward the optimal hyperparameter configuration at an exponential rate, provided the selection pressure is sufficiently large. The result makes the roles of noise strength, population diversity, and selection pressure explicit and quantitative.
  • Algorithms and experiments (Section 5). We propose a reduced training algorithm (Algorithm 2) in which the inner SGD loop is replaced by sampling from, or time-averaging over, the parameter equilibrium. Experiments on quadratic and Himmelblau bilevel problems, and on a CartPole deep reinforcement learning task, validate the large-population limit, the two-time-scale regime, and the practical benefit of effective fitness estimation.
The remainder of the paper is organised as follows. In Section 2, we introduce the stochastic particle model describing a population of neural networks undergoing both training and evolutionary dynamics. We also derive the corresponding kinetic model in the many-agent limit $N\to\infty$, first formally and then rigorously. In Section 3, we analyse the two-time-scale limit, leading to a reduced macroscopic selection–mutation dynamics for hyperparameters and its connection with bilevel optimisation and replicator–mutator models. In Section 4, we study this reduced dynamics and identify assumptions under which the population mean moves toward the fittest hyperparameter configuration. Numerical examples are presented in Section 5, and concluding remarks and perspectives are discussed in Section 6.
2 Mathematical modelling of population-based training
2.1 The collective training mechanism
Population-Based Training (PBT) aims to find the best Neural Network (NN) for a given task by optimising both the set of parameters $\theta$ and of hyperparameters $h$ which we assume take values in $\Theta\subseteq\mathbb{R}^{d_{\theta}},H\subseteq\mathbb{R}^{d_{h}}$ respectively. The number of parameters is typically of several order of magnitude larger than the number of hyperparameters, $d_{\theta}\gg d_{h}$.
PBT employs a set of $\in\mathbb{N} NN$s, which we will also call agents
$(\theta^{i},h^{i})\in\Theta\times H\qquad i=1,\dots,N$
that are iteratively updated according to two mechanisms:
  • Loss $\mathcal{L}$ minimisation via independent learning dynamics;
  • Fitness $\mathcal{F}$ maximisation via collective interaction.
Parameters training.
The loss minimisation problem is solved by gradient-based algorithms such as SGD, RMSprop or ADAM-type optimisers. While the learning dynamics is used only to update the parameters $\theta^{i}$, it typically depends on the hyper-parameters $h^{i}$ since they could include, for instance, the learning step, regularisation or noise strength. To stress this dependence we consider a loss $\mathcal{L}=\mathcal{L}(\theta,h)$ which depends on both $\theta$ and $h$.
In the case of SGD training, the parameter dynamics can be approximated by a stochastic differential equation (SDE) of the form
$d\theta^{i}=-\nabla_{\theta}\mathcal{L}(\theta^{i},h^{i})\,dt+\sqrt{2\beta^{-1}(h^{i})}\Sigma^{1/2}(\theta^{i})\,dB^{i}_{t},$ (1)
where $(B_{t}^{i})_{t\geq 0}$ are independent $d_{\theta}$-dimensional Brownian motions. The matrix $\Sigma(\theta)$ is positive definite and captures the variance induced by stochastic loss sampling, while $\beta(h)>0$ determines the noise strength, and typically depends on the hyperparameters via the learning rate (Cheng et al., 2020; Li et al., 2019b, 2021). RMSprop and Adam adaptive training algorithms can be also modelled via SDE of type, by adding auxiliary random variables into the dynamics (Malladi et al., 2022), see Remark 1 for more details.
Hyperparameters evolution.
The fitness function $\mathcal{F}$ is conceptually different from the loss $\mathcal{L}$, as it is not directly related to the training dynamics. For instance, while $\mathcal{L}$ may depend on the training data, $\mathcal{F}$ typically evaluates the generalisation performance of an agent-NN on a validation dataset. Moreover, the fitness function does not need to be differentiable. In general, it depends on both variables: $\mathcal{F}=\mathcal{F}(\theta,h)$.
The fitness maximisation is addressed by PBT via genetic-type dynamics where the best performing agents (that is, the fittest) are copied with a certain rate. To keep the number of agents limited, the worst performing ones are discarded, or they are discarded randomly. At a slower rate (with respect to the inner learning dynamics) each agent $(\theta,h)$ is subject to a genetic update of type
$\begin{split}\theta^{\prime}&=\theta_{*}\\ h^{\prime}&=h_{*}+\sigma\xi\end{split}$ (2)
where $(h_{*},\theta_{*})$ is an agent sampled among the best ones in the following way
$\mathbb{P}\left((\theta_{*},h_{*})=(\theta^{i},h^{i})\right)=\frac{\exp(\mathcal{F}(\theta^{i},h^{i}))}{\sum_{j=1}^{N}\exp(\mathcal{F}(\theta^{j},h^{j}))}\,.$
Note that the parameter $\theta$ is copied to $\theta_{*}$, and mutation acts only at the hyper-parameter level via a mutation vector $\xi\sim\mathcal{N}(0,I_{d})$ with strength $\sigma>0$. If the search space $H$ is bounded and convex, an additional projection is added to ensure $h^{i}\in H$ at all times, by modifying $h^{\prime}$ in (2) as $h^{\prime}\leftarrow\textup{Project}_{H}(h_{*}+\sigma\xi)$.
A semi-discrete time evolution.
The resulting algorithm can be seen as a continuum training of the parameters $\theta^{i}$, interrupted by genetic updates (or jumps) where $(\theta^{i},h^{i})$ are resampled among the fittest agents. See Algorithm 1 for a high-level description of the PBT evolution.
In Algorithm 1 the parameter $\tau\in(0,1)$ determines the frequency of such evolutionary interruptions. For $\tau=1$ we have that all agents perform a genetic update simultaneously. For $\tau\ll 1$, interruptions are more frequent, but agents are less likely to undergo resampling. Note that for any $\tau\in(0,1]$ the expected total number of jumps per agent per unit time is always 1.
Algorithm 1 Population-Based Training
Initialize $\in\mathbb{N} NN$s with parameters $\theta^{i}\in\Theta$ and hyperparameters $h^{i}\in H$, $i=1,\dots,N$
Set $\tau\in(0,1]$$\sigma>0$, $t_{\max}>0$
$t\leftarrow 0$
while $t<t_{\max}$ do
  for $i=1,\dots,N$ do
   Train $\theta^{i}$ via SGD (1) for $t,t+\tau) \triangleright$ parallel NN training
   $F^{i}\leftarrow\mathcal{F}(\theta^{i},h^{i})$ $\triangleright$ evaluate fitness
  end for
  for $i=1,\dots,N$ do $\triangleright$ population interaction
   with probability $\tau$:
     select a NN $(\theta^{j},h^{j})$ proportionally to $\exp(F^{j})$
     sample random mutation vector $\xi$
     $\tilde{\theta}^{i}\leftarrow\theta^{j}$ $\triangleright$ copy parameter
     $\tilde{h}^{i}\leftarrow h^{j}+\sigma\xi$ $\triangleright$ copy and perturb hyperparameter
     $\tilde{h}^{i}\leftarrow\textup{Project}_{H}(\tilde{h}^{i})$ $\triangleright$ project to search space
  end for
  $\theta^{i}\leftarrow\tilde{\theta}^{i},h^{i}\leftarrow\tilde{h}^{i}$ for all $i=1,\dots,N$
  $t\leftarrow t+\tau$
end while
2.2 Description in the many-agent limit
When dealing with a large interacting system of $N\gg 1$ agents $(\theta^{i},h^{i}),i=1,\dots,N$, it is convenient to provide a statistical description of the system. Rather than tracking the microscopic trajectories of each agent $(\theta^{i},h^{i})$, we aim to derive a model which quantifies the probability of finding a certain agent with features $(\theta,h)$ at a given time $t\geq 0$. The model can be formally obtained by taking the many agent limit $N\to\infty$.
We consider a probability distribution $f=f(t,\theta,h)$ in the parameter/hyperparameter phase space $\Theta\times H\subseteq\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{h}}$ which we assume it well approximates the empirical distribution
$f(t,\theta,h)\;\approx\;f^{N}(t,\theta,h)=\frac{1}{N}\sum_{i=1}^{N}\delta(\theta-\theta^{i}(t))\delta(h-h^{i}(t))\qquad\textup{for}\quad N\gg 1\,$, (3)
where $\delta(\cdot)$ is the Dirac delta distribution centred at 0.
Training operator.
When a random variable $\theta^{i}$ evolves according to a SDE of type (1), the evolution of its probability law is determined by a Fokker–Planck operator. In (1), we note that only the parameters $\theta^{i}$ evolve, while the hyperparameters $h^{i}$ are fixed. Therefore, for $f=f(t,\theta,h)$, the Fokker–Planck operator associated with the training dynamics is given by
$T_{\mathcal{L}}[f]:=\nabla_{\theta}\cdot\bigl(\nabla_{\theta}\mathcal{L}(\theta,h)\,f\bigr)+\beta^{-1}(h)\nabla_{\theta}^{2}:(\Sigma(\theta)\,f)\,$, (4)
see, for instance, (Bogachev et al., 2022, Chapter 6). Above, $A:B=\mathrm{tr}(A^{\top}B)$ is the usual Frobenius inner product. For Adam-based training, a similar Fokker–Planck operator can be also obtained via the corresponding SDE formulation, see Remark 1.
Evolutionary operator.
It is during the collective genetic-type jump that agents interact. Assumption (3) can also be understood as a so-called propagation of chaos assumption on the agents (Sznitman, 1991). This means that for large population sizes the correlation between the agents become weaker and weaker, until they become independent as $N\to\infty$. As agents are indistinguishable, this also implies that they are identically distributed with law $f$.
If a genetic update occurs at time $t$, the probability density of copying an agent with state $(\theta,h)$ is then given by
$g(t,\theta,h)=\frac{e^{\mathcal{F}(\theta,h)}\,f(t,\theta,h)}{\iint e^{\mathcal{F}(\theta^{\prime},h^{\prime})}\,f(t,\theta^{\prime},h^{\prime})\,d\theta^{\prime}\,dh^{\prime}}\,.$
Instead, the hyperparameter mutation in (2) corresponds to a convolution with a probability kernel $K$ in the $h$ variable, namely
$(K_{\sigma}*_{h}g)(t,\theta,h)=\int g(t,\theta,h+\sigma\xi)\,K(d\xi)\,.$
The selection–mutation mechanism outlined in Algorithm 1 therefore corresponds, at the distribution level, to the reweighting–convolution operator
$E_{\mathcal{F}}[f]:=K_{\sigma}*_{h}\left(\frac{e^{\mathcal{F}}f}{\int e^{\mathcal{F}}df}\right)-f\,.$ (5)
The term f-f is required because a new agent is created due to the jump, while the pre-jump agent is deleted in the process. This ensures that the total amount of agents does not change during the computation.
A PDE model.
As discussed earlier, the parameter $\tau\in(0,1]$ describes how frequently each agent has the possibility of undergoing a jump. It can be interpreted as a time-discretisation parameter for the collective evolutionary dynamics where, in the limit $\tau\to 0$, jumps may occur at any time and asynchronously. As $N\to\infty$ and $\tau\to 0$, we formally obtain a continuum description of Algorithm 1, in which the time evolution of the kinetic distribution $f$ is determined by both the training and evolutionary operators
$\frac{\partial f}{\partial t}=T_{\mathcal{L}}[f]+E_{\mathcal{F}}[f]\,.$ (6)
Remark 1
The training operator for adaptive algorithms such as RMSprop or Adam can be equivalently derived from their continuous SDE approximations, as for instance in (Malladi et al., 2022). The main difference with respect to SGD is that the parameter state space must be augmented to account for the momentum variable $m$, or for other auxiliary variables appearing in the dynamics.
For instance, in the case of Adam the SDE reads
$\begin{cases}d\theta^{i}=-P^{-1}(h^{i},u^{i})m^{i}\,dt,\\ dm^{i}=\alpha(h^{i})\bigl(\nabla_{\theta}\mathcal{L}(\theta^{i},h^{i})-m^{i}\bigr)\,dt+\sqrt{2\beta^{-1}(h^{i})}\,\Sigma^{1/2}(\theta^{i})\,dB^{i},\\ du^{i}=\gamma(h^{i})\bigl(\operatorname{diag}(\Sigma(\theta^{i}))-u^{i}\bigr)\,dt,\end{cases}$
where $\alpha(h)$ and $\gamma(h)$ are hyperparameters and $P(h,u)$ is a preconditioning matrix. At the PDE level, the associated distribution must also be extended and takes the form $f=f(t,\theta,h,m,u)$. As will be discussed in Section 3, these auxiliary variables become irrelevant in the two-time-scale limit, simplifying the description.
Remark 2
A related line of work studies the optimal design of training protocols by casting reduced learning dynamics as an optimal control problem, often in simplified teacher–student models where the dynamics closes on a few order parameters (yielding low-dimensional ODEs (Mignacco and Mori, 2025)). Our approach is complementary in scope and object: rather than controlling a single reduced trajectory, we derive an averaged population dynamics for the hyperparameter distribution $\rho(t,h)$, obtained from the large population limit and a two-time-scale (fast training/slow selection–mutation) limit. This averaged population equation can be regarded as a state equation for future protocol-design problems, including but not limited to optimal control.
2.3 Propagation of chaos and convergence
We show now that the connection between Algorithm 1 and the kinetic model (6) can be made rigorous. We will consider weak measure solutions to (6), see Definition 12 for a precise characterisation. We restrict to analysis to the case of SGD parameters training with isotropic noise.
Assumption 3
  1. 1.
    Parameters are trained with SGD via (1) with $\beta>0$ fixed and $\Sigma(\theta)\equiv I$.
  2. 2.
    The loss $\mathcal{L}\in C^{1}(\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{h}})$ is globally Lipschitz continuous.
  3. 3.
    The fitness $\mathcal{F}\in C(\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{h}})$ is globally Lipschitz continuous and bounded with $0<\inf{\mathcal{F}}<\sup\mathcal{F}<\infty$ .
To compare the empirical distribution $f^{N,\tau}$ (we stress now also the dependence on $\tau\in(0,1)$) and the kinetic distribution $f$, we will employ the Bounded-Lipschitz (BL) norm between measures, also known as flat, or Dudley metric (Dudley, 2018). For any Borel measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ it is defined as
$\|\mu\|_{{\mathrm{BL}}}:=\sup\left\{\int\phi\,d\mu\;\middle|\;\|\phi\|_{\infty}<1/2,\|\phi\|_{\textup{Lip}}<1\right\}$ (7)
where $\|\phi\|_{\infty}:=\sup_{x}|\phi(x)|$ and $\|\phi\|_{\textup{Lip}}:=\sup_{x\neq y}|\phi(x)-\phi(y)|/|x-y|$.
Theorem 4
Let Assumption 3 hold, $t_{\max}>0$ be a computational time horizon, and $f_{0}\in\mathcal{P}_{q}(\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{h}}),q>1$ an initial distribution. Consider the multi-agent system generated by Algorithm 1 $\{(\theta^{i}(t),h^{i}(t))_{t\in[0,t_{\max}]}\}_{i=1}^{N}$ with $f^{N,\tau}(t)$ being the associated empirical measure and $(\theta^{i}(0),h^{i}(0))$ being independent and $f_{0}$-distributed.
Then, there exists a weak measure solution $f\in C([0,t_{\max}],\mathcal{P}_{q}(\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{h}}))$ to (6) with initial data $f_{0}$, and it holds
$\mathbb{E}\|f^{N,\tau}(t)-f(t)\|_{\mathrm{BL}}\longrightarrow 0\qquad\textup{as}\quad N\to\infty,\;\tau\to 0$
for all $t\in[0,t_{\max}]$.
Since $f(t)$ is deterministic, Theorem 4 implies that the stochastic multi-agent system becomes deterministic in the limit $N\to\infty$. Moreover, the result validates the ansatz $f^{N}\approx f$ (3) used in the derivation of the kinetic PDE (6). We prove the theorem by first considering the limit $N\to\infty$, via a coupling technique and the optimal transport formulation of the BL norm. Then, the limit $\tau\to 0$ is considered by exploiting the stability both the training and evolutionary operators with respect to the BL norm. We refer to Appendix B for the details.
2.4 Modelling extensions
In genetic algorithms, many variants of selection procedures have been proposed, see (Jebari, 2013) for a review. In more sophisticated mechanisms, poorly performing agents are more likely to undergo a genetic update, or an agent’s quality is determined by its rank within the population rather than by its fitness value.
Substitution of NNs performing poorly.
In Algorithm 1, every NN is equally likely to be updated, with probability $\tau$ at every time $t_{n}=n\tau$. The genetic update frequency could also be made fitness-dependent by prescribing a higher replacement probability for agents that have poor fitness values as done in (Jaderberg et al., 2017). This amounts, at the PDE level, to modifying the loss term in (5). In this case, agents are no longer selected uniformly for removal, but rather with a bias favouring the least fit ones. A natural choice is to use the same exponential weighting as in the gain term, but applied to the negative fitness $\mathcal{F}(\theta,h)$. This more elaborate algorithmic strategy would lead to the operator
$\widehat{E}_{\mathcal{F}}[f]:=K_{\sigma}*_{h}\left(\frac{e^{\mathcal{F}}f}{\int e^{\mathcal{F}}df}\right)-\frac{e^{-\mathcal{F}}f}{\int e^{-\mathcal{F}}df}.$
Rank-based selection.
In Algorithm 1, the probability of a NN agent $(\theta^{i},h^{i})$ to be selected for duplication is proportional to its fitness $\mathcal{F}(\theta^{i},h^{i})$. Another common strategy consists of picking agents with respect to their rank among the ensemble. For instance, one can decide that only the top $20\%$ NN are copied when performing a genetic update, see (Jaderberg et al., 2017).
Mathematically, given a group of NNs $(\bm{\theta},\bm{h})=\{(\theta^{i},h^{i})\}_{i=1}^{N}$, the function rank is defined as
$\mathrm{Rank}(\theta^{i},h^{i}\,|\,\bm{\theta},\bm{h}):=\#\{j\,|\,\mathcal{F}(\theta^{j},h^{j})\geq\mathcal{F}(\theta^{i},h^{i})\}\,.$
This can be extended for an arbitrary probability distribution $f\in\mathcal{P}(\mathbb{R}^{d_{\theta}},\mathbb{R}^{d_{h}})$ as
[…]
Therefore, at least formally, it is possible to derive kinetic description of the collective training dynamics where rank-based selection is used, as done in Section 2.2 for fitness-based selection. A rigorous convergence analysis is more delicate, and we leave it for future work. In the literature, rank-based particle interactions have been considered in (Bencheikh and Jourdain, 2022; Demircigil and Tomasevic, 2025).
3 Two-time-scale limit
Fast training, slow evolution.
The two processes, loss minimisation and fitness maximisation, happen at two different time scales. Typically, several training steps are performed (of order $10^{4}$ (Jaderberg et al., 2017), for instance) before a single evolutionary update. This also reflects the dimensionality unbalance between the number of parameters and of hyperparameters. To balance the two processes we introduce a scaling parameter $\varepsilon>0$ and study the rescaled model
$\frac{\partial f}{\partial t}=\frac{1}{\varepsilon}T_{\mathcal{L}}[f]+E_{\mathcal{F}}[f]\,.$ (8)
At the algorithmic level, this corresponds to rescaling the time of the parameter optimisation (1), making the loss training longer of a factor $1/\varepsilon$.
Limit as $\varepsilon\to 0$.
We are interested in studying the evolution of the hyperparameter distribution
$\rho(t,h)=\int f(t,\theta,h)d\theta$
by averaging out the fast time evolution for $\varepsilon\ll 1$. In the limit $\varepsilon\to 0$, equation (8) reduces to
$T_{\mathcal{L}}[f]=0\,.$
which implies that $f$ should be at equilibrium with respect to $T_{\mathcal{L}}$, taking the form
$f(t,h,\theta)=\rho(t,h)\,\mu^{\infty}_{\mathcal{L}}(\theta|h)\,.$ (9)
with $T_{\mathcal{L}}[\mu^{\infty}(\cdot|h)]=0$. This means that, given a certain hyperparameter $h\in H$, the distribution on the space of parameters is fixed in time, and given by the conditional probability $\mu^{\infty}_{\mathcal{L}}(\theta|h)$. At this slow time scale, only the hyperparameters distribution $\rho=\rho(t,h)$ evolves.
The specific form of the equilibria $\mu^{\infty}_{\mathcal{L}}$ depends on the training dynamics used. Since $T_{\mathcal{L}}$ corresponds to the Fokker–Planck operator, they typically take the form of exponential distributions associated with a given potential. For instance, when considering SGD with isotropic noise ($\Sigma\equiv I_{d}$ in (1)), the equilibrium is given by a Gibbs distribution associated with the loss function:
$\mu^{\infty}_{\mathcal{L}}(\theta|h)=\frac{e^{-\beta(h)\mathcal{L}(\theta,h)}}{\int e^{-\beta(h)\mathcal{L}(\theta,h)}d\theta}\,$, (10)
see, for instance, (Pavliotis, 2014, Proposition 4.6). Existence of an invariant distribution and convergence towards it is related for Fokker–Planck operator in more general settings have been investigated in (Ottobre and Pavliotis, 2011; Achleitner et al., 2015; Mandt et al., 2017). Also, they might not have an explicit form due to the possible noise degeneracy. Specific results for SGD, RMSProp, and Adam-type algorithms have been obtained in (Mandt et al., 2017; Suzuki, 2023).
Under ansatz (9), if we test the solution $f$ against $\phi=\phi(h)$ we obtain $\langle\phi,T_{\mathcal{L}}[f]\rangle=0$ and
$\langle\phi,E_{\mathcal{F}}[f]\rangle$ $=\iiint\left(\frac{e^{\mathcal{F}(\theta,h)}}{\int e^{\mathcal{F}}\,df}\,\phi(h+\sigma\xi)-\phi(h)\right)K(\xi)\,\rho(t,h)\,\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\xi\,d\theta\,dh$
$=\iint\left(\int\frac{e^{\mathcal{F}(\theta,h)}}{\int e^{\mathcal{F}}\,df}\,\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\right)\phi(h+\sigma\xi)\,K(\xi)\,\rho(t,h)\,d\xi\,dh\;-\;\int\phi(h)\,\rho(t,h)\,dh.$
Averaged equation.
Let $\overline{\mathcal{F}}=\overline{\mathcal{F}}(h)$ be what we call the effective Gibbs-averaged fitness
$\overline{\mathcal{F}}(h):=\log\int e^{\mathcal{F}(\theta,h)}\mu^{\infty}_{\mathcal{L}}(\theta|h)d\theta\,.$ (11)
Note that the normalizing constant remains unchanged
$\int e^{\mathcal{F}}\,df=\iint e^{\mathcal{F}(\theta,h)}\,\rho(t,h)\,\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\,dh=\int\exp\!\left(\log\int e^{\mathcal{F}(\theta,h)}\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\right)\rho(t,h)\,dh=\int e^{\overline{\mathcal{F}}}\,d\rho\,$,
leading to
$\langle\phi,E_{\mathcal{F}}[f]\rangle=\langle\phi,\overline{E}_{\overline{\mathcal{F}}}[\rho]\rangle:=\iint\left(\frac{e^{\overline{\mathcal{F}}(h)}}{\int e^{\overline{\mathcal{F}}}\,d\rho}\,\phi(h+\sigma\xi)-\phi(h)\right)K(\xi)\,\rho(t,h)\,d\xi\,dh.$ (12)
Altogether, we obtain in the limit $\varepsilon\to 0$ the averaged equation for $\rho=\rho(t,h)$
$\frac{\partial\rho}{\partial t}=\overline{E}_{\overline{\mathcal{F}}}[\rho]\,.$ (13)
Therefore, in the two-time scale limit the dynamics reduces to a pure selection-mutation evolution for the effective fitness $\overline{\mathcal{F}}$. Equations of this type have been studied in (Del Moral and Miclo, 2000; Del Moral, 2004) to study the many agent limit of genetic-type algorithms via Feynman–Kac distribution flows.
Remark 5
The key point in the derivation of the averaged PDE (13) is the equilibrium ansatz (9). The emergence of equilibrium in parameter space depends only on the scale separation between training and evolution, and its precise form depends on the specific algorithm used for parameter training. Therefore, the type of evolutionary strategy employed does not affect the emergence of the lower-dimensional PDE. The same reduction can thus be carried out, for instance, for rank-based selection or for the other extensions discussed in Section 2.4.
Remark 6
For stochastic optimisation algorithms, the long-time behaviour may exhibit non-equilibrium features, including persistent probability currents (Chaudhari and Soatto, 2018; Li et al., 2022). In (Li et al., 2022), the authors relate this behaviour to the implicit regularisation properties of SGD by showing the existence of slow dynamics along manifolds of minimizers. In our modelling framework, this corresponds to having an implicit dependence in the Fokker–Planck operator $T_{\mathcal{L}}$ by $\varepsilon$. In this case, more accurate macroscopic models can de derived by performing a Chapman–Enskog-type expansion (Bobylev, 2020) of the training operator.
3.1 Connections to bi-level optimisation problems
We shed some light on the role of the effective fitness $\overline{\mathcal{F}}$ by relating it to bi-level optimisation problems. The NN training, indeed, can be cast as a problem of type
$\begin{split}h^{\star}&\in\underset{h\in H}{\arg\!\max}\,\mathcal{F}(\theta^{\star}(h),h)\\ &\textup{subject to}\quad\theta^{\star}(h)\in\underset{\theta\in\Theta}{\textup{argmin}}\,\mathcal{L}(\theta,h)\end{split}$ (14)
where one sequentially aims to minimize the loss $\mathcal{L}$ and maximize the fitness $\mathcal{F}$ by tuning parameters and hyperparameters. For each fixed $h$, $\overline{\mathcal{F}}(h)$ fits into the variational–inference viewpoint: the equilibrium measure $\mu^{\infty}_{\mathcal{L}}$ acts as a prior, possibly concentrating on minimisers of $\mathcal{L}(\cdot,h)$, while $\mathcal{F}(\theta,h)$ plays the role of a log-likelihood.
A penalisation viewpoint.
Now, let us consider the case where the equilibrium is explicitly given by the Gibbs measure (10), with fixed $\beta>0$. We stress the dependence on $\beta$ by writing the averaged loss as $\overline{\mathcal{F}}_{\beta}$. Let $\mathcal{H}$ be the log-entropy, we rewrite the KL divergence as
$\textup{KL}(\nu\,|\,\mu_{\mathcal{L}}^{\infty})$ $=\int\nu(\theta)\log(\nu(\theta))d\theta-\int\nu(\theta)\log\left(\frac{e^{-\beta\mathcal{L}(\theta,h)}}{\int e^{-\beta\mathcal{L}}d\theta}\right)d\theta$
$=\mathcal{H}[\nu]+\beta\mathbb{E}_{\nu}[\mathcal{L}(\theta,h)]+\log\int e^{-\beta\mathcal{L}}d\theta$
Let us denote $\overline{\mathcal{L}}_{\beta}(h):=-(1/\beta)\log\int e^{-\beta\mathcal{L}(\theta,h)}d\theta$ for which, by the Laplace principle (Dembo and Zeitouni, 2010), it holds
$\overline{\mathcal{L}}_{\beta}(h)\longrightarrow\inf_{\theta}\mathcal{L}(\theta,h)\qquad\textup{as}\quad\beta\to\infty\,.$
We obtain a further characterisation of the effective fitness as
$\overline{\mathcal{F}}_{\beta}(h)=\sup_{\nu}\Big(\mathbb{E}_{\nu}\left[\mathcal{F}(\theta,h)-\beta\left(\mathcal{L}(\theta,h)-\overline{\mathcal{L}}_{\beta}(h)\right)\right]-\mathcal{H}[\nu]\Big)\,.$
which has a clear interpretation in terms of the bi-level optimisation problem (14). Indeed, in view of (3.1), we can see $\overline{\mathcal{F}}_{\beta}(h)$ as an entropy regularized version of a penalized fitness
${\mathcal{P}}_{\beta}(h):=\mathcal{F}(\theta,h)-\beta\left(\mathcal{L}(\theta,h)-\inf_{\theta}\mathcal{L}(\theta,h)\right)$
where $\beta$ here plays the role of the penalisation parameter. This is a standard approach in solving the bi-level optimisation problem, where (14) is viewed as a constrained optimisation problem, see, for instance, (Kwon et al., 2023), and for which it holds $\max_{h}{\mathcal{P}}_{\beta}(h)\to\max_{h}\mathcal{F}(h,\theta^{\star}(h))$ as $\beta\to\infty$. The following lemma quantifies this intuition and show that the same holds for the averaged problem $\max_{h}\overline{\mathcal{F}}_{\beta}(h)$. See Appendix C.2 for the proof.
Assumption 7
Assume the hyperparameter space $H\subset\mathbb{R}^{d_{h}}$ to be compact and the loss function $\mathcal{L}=\mathcal{L}(\theta,h)$ to satisfy:
  1. 1.
    For every $h\in H$, $\mathcal{L}(\cdot,h)\in C^{4}(\mathbb{R}^{d_{\theta}})$,
  2. 2.
    For every $h\in H$, the map $\theta\mapsto\mathcal{L}(\theta,h)$ admits a unique global minimizer $\theta^{\star}(h)\in\mathbb{R}^{d_{\theta}}$ at which the Hessian is positive definite: there exists $\lambda>0$ such that
    $\nabla_{\theta}^{2}\mathcal{L}(\theta^{\star}(h),h)\succeq\lambda I_{d_{\theta}}\qquad\textup{for all}\quad h\in H\,.$
  3. 3.
    For every $r>0$,
    $\inf\Bigl\{\mathcal{L}(\theta,h)-\mathcal{L}(\theta^{\star}(h),h):h\in H,\ |\theta-\theta^{\star}(h)|\geq r\Bigr\}>0.$
  4. 4.
    The Gibbs weights have uniformly integrable first moments:
    $\sup_{h\in H}\int_{\mathbb{R}^{d_{\theta}}}|\theta|\,\exp\bigl(-(\mathcal{L}(\theta,h)-\mathcal{L}(\theta^{\star}(h),h))\bigr)\,d\theta<\infty.$
Lemma 8
Let $\mathcal{F},\mathcal{L}$ satisfy Assumption 3 and 7. Then, for all $\beta\geq 1$ the Gibbs-averaged fitness (3.1) satisfies
$|\overline{\mathcal{F}}_{\beta}(h)-\mathcal{F}(\theta^{\star}(h),h)|\lesssim\frac{1}{\sqrt{\beta}}\qquad\textup{for all}\quad h\in H.$
The proof is based on the quantitative Laplace principle derived in (Hasenpflug et al., 2024), where the authors provide a rate for the limit $\mu^{\infty}_{\mathcal{L}}(\cdot|h)\to\delta_{\theta^{\star}(h)}$, as $\beta\to\infty$, in terms on Wasserstein distances. Recall $\beta>0$ is the inverse of the temperature in the Gibbs measures $\mu^{\infty}_{\mathcal{L}}$. For simplicity, we applied the result in the case where $\mathcal{L}(\cdot,h)$ attains a unique global minimum for each $h\in H$, but the assumptions can be made more general to include manifolds of solutions, see (Hasenpflug et al., 2024) for more details. Assumption 7 (4) ensures that the Gibbs equilibria are well-defined, while (3) ensures that there is no diverging sequence of local minima $\theta^{k}$ such that $\mathcal{L}(\theta^{k},h)\to\min\mathcal{L}(\cdot,h)$ as $k\to\infty$.
3.2 Diffusion–reaction approximation in the slow mutation regime
To relate (13) with the evolutionary PDEs typically used for biological systems (Perthame, 2007; Ackleh et al., 2016; Alfaro and Veruete, 2019), we consider the following re-scaling for $\nu>0$
$t\to t/\nu\,,\quad\sigma\to\sqrt{\nu}\sigma\,,\quad\overline{\mathcal{F}}\to\nu\overline{\mathcal{F}}\,.$ (15)
Slow mutation limit.
Under this rescaling of the fitness $\overline{\mathcal{F}}$, we obtain as $\nu\to 0$
$e^{\nu\overline{\mathcal{F}}}=1+\nu\overline{\mathcal{F}}+\mathcal{O}(\nu^{2})$
and, by integrating, also $\int e^{\nu\overline{\mathcal{F}}}d\rho=1+\nu\int\overline{\mathcal{F}}d\rho+o(\nu)\,.$ Using the formula: $(1+\nu x)/(1+\nu y)=1+\nu(x-y)+o(\nu)$ we get that the exponential weights can be approximated as
$\frac{e^{\nu\overline{\mathcal{F}}}}{\int e^{\nu\overline{\mathcal{F}}}d\rho}=1+\nu\left(\overline{\mathcal{F}}-\langle\overline{\mathcal{F}},\rho\rangle\right)+o(\nu)\,.$
The Taylor expansion for a smooth test function $\phi=\phi(h)$ instead gets us
$\int\phi(h^{\prime})K(d\xi)$ $=\int\phi(h+\sqrt{\nu}\sigma\xi,\theta)K(d\xi)$
$=\int\left(\phi(h)+\sqrt{\nu}\sigma\xi\cdot\nabla\phi(h)+\frac{1}{2}\nu\sigma^{2}\xi^{\top}\nabla^{2}\phi(h)\xi\right)K(d\xi)$
$=\phi(h)+\nu\frac{\sigma^{2}}{2}\Delta\phi(h)+o(\nu)\,.$
where we used that $\mathbb{E}[\xi]=0$, $\mathbb{E}[\xi\tilde{\xi}^{\top}]=I$ for $\xi,\tilde{\xi}\sim K$. Altogether we obtain
$\frac{e^{\nu\overline{\mathcal{F}}(h)}}{Z_{\nu\overline{\mathcal{F}}}[\rho]}\int\phi(h+\sigma\xi)K(d\xi)$ $=(1+\nu(\overline{\mathcal{F}}(h)-\langle\overline{\mathcal{F}},\rho\rangle)\left(\phi(h)+\nu\frac{\sigma^{2}}{2}\Delta\phi\right)+o(\nu)$
$=\phi(h)-\nu\left(\overline{\mathcal{F}}(h)-\langle\overline{\mathcal{F}},\rho\rangle\right)\phi(h)+\nu\frac{\sigma^{2}}{2}\Delta\phi(h)+o(\nu)\,$,
which leads, in turn, to
$\langle\phi,\overline{E}^{\nu}_{\overline{\mathcal{F}}}[\rho]\rangle=\left\langle K_{\sqrt{\nu}\sigma}\ast\phi,\frac{e^{\nu\overline{\mathcal{F}}}}{Z_{\nu\overline{\mathcal{F}}}[\rho]}\rho\right\rangle-\langle\phi,\rho\rangle=\nu\left(\left\langle\phi,\overline{\mathcal{F}}\rho\rangle-\langle\overline{\mathcal{F}},\rho\rangle\langle\phi,\rho\right\rangle\right)+\nu\frac{\sigma^{2}}{2}\langle\Delta\phi,\rho\rangle+o(\nu)\,.$
A replicator-mutation PDE.
Therefore, (13) under the slow selection rescaling (15) formally leads in the limit $\nu\to 0$ to a diffusion-reaction equation
$\frac{\partial\rho}{\partial t}=\rho\left(\overline{\mathcal{F}}(h)-\langle\overline{\mathcal{F}},\rho\rangle\right)+\frac{\sigma^{2}}{2}\Delta\rho\,$, (16)
which has been largely studied in the context of evolutionary equations in mathematical biology (Perthame, 2007). The mutation mechanism is now described by a Laplacian term, while selection is described by a source term where particles at $h$ are created with rate $\overline{\mathcal{F}}(h)$ destroyed uniformly with a uniform rate $\langle\overline{\mathcal{F}},\rho\rangle$. Note that the terms are balanced so that the total mass remains constant during the evolution. Equations of type (16) are known in the sampling literature also as birth-death or replicator dynamics (Lu et al., 2023), and can be cast as gradient flows with respect to the Hellinger–Kantorovich (Wasserstein–Fisher–Rao) geometry when additional drift terms are present (Liero et al., 2018; Gallouët et al., 2019). The slow mutation limit therefore reveals a remarkable connection between genetic-type algorithms and gradient flows.
4 Mean convergence toward the fittest solution
In this section we study the evolution of the averaged PDE (13) to get an understanding on the hyperparameters evolution at large times. Existence of measure solutions to (13) have been studied in (Del Moral and Miclo, 2000; Del Moral, 2004; Borghi, 2026). In (Del Moral and Miclo, 2002) the authors show that averaged evolutionary operator (12) is stable, and solutions to (13) exponentially converge to the unique steady state, under suitable assumptions. This is achieved by using semigroup techniques and Dobrushin’s ergodic coefficient, but an explicit characterisation of the steady states is missing.
Here, we are interested here in understanding whether we can expect the hyperparameters to converge towards the fittest solution with respect to the averaged objective
$h^{\star}\in\underset{h\in\mathbb{R}^{d_{h}}}{\mathrm{argmax}}\;\overline{\mathcal{F}}(h)\,.$ (17)
To do so, we propose an analysis based on bounds on the moments evolutions, and a quantitative version of the Laplace principle.
4.1 Moments evolution
Denote with $\textup{m}(\rho)$ the mean of $\rho\in\mathcal{P}(\mathbb{R}^{d_{h}})$, $\textup{m}(\rho):=\int h\,\rho(dh)\,$, and the particle energy, or second moment, as $\mathrm{En}(\rho):=(1/2)\int|h|^{2}\rho(dh)\,.$ For notational convenience we also define the action of reweighing the agents distribution $\rho$ with respect to the average fitness $\overline{\mathcal{F}}$ as
$G_{\overline{\mathcal{F}}}[\rho](dh):=\frac{e^{\overline{\mathcal{F}}(h)}\rho(dh)}{\int e^{\overline{\mathcal{F}}(h^{\prime})}\rho(dh^{\prime})}\,.$ (18)
We note that the evolutionary operator (12) can be compactly be written as $\overline{E}_{\overline{\mathcal{F}}}[\rho]=K_{\sigma}\ast G_{\overline{\mathcal{F}}}[\rho]$.
Along solutions $\rho=\rho(t,h)$ of the averaged PDE (13), it holds
$\frac{d}{dt}\textup{m}(\rho)=\textup{m}\left(G_{\overline{\mathcal{F}}}[\rho]\right)-\textup{m}(\rho)\,.$ (19)
This can be shown by directly applying the weak formulation (12) of the evolutionary operator. Since the mutation vectors are mean-zero, mutation does not affect the mean evolution, which is fully determined by the selection mechanism.
The evolution of the energy $\mathrm{En}(\rho)$ is instead given by
$\frac{d}{dt}\mathrm{En}(\rho(t))$ $=\iint|h+\sigma\xi|^{2}G_{\overline{\mathcal{F}}}[\rho(t)](dh)-\mathrm{En}(\rho(t))$
$=\mathrm{En}\left(G_{\overline{\mathcal{F}}}[\rho(t)]\right)-\mathrm{En}(\rho(t))+d_{h}\sigma^{2}\,$,
where we used that $\int|\xi|^{2}K(d\xi)=d_{h}$ since $K$ is the standard $d_{h}$-dimensional Gaussian measure. From the moments’ evolution, then, we understand that any equilibrium $\rho^{\infty}$ should satisfy the equations
$\begin{split}\textup{m}(\rho^{\infty})&=\textup{m}\left(G_{\overline{\mathcal{F}}}[\rho^{\infty}]\right)\\ \mathrm{En}(\rho^{\infty})&=\mathrm{En}\left(G_{\overline{\mathcal{F}}}[\rho^{\infty}]\right)+d_{h}\sigma^{2}\,.\end{split}$ (20)
It is also interesting to note that there is a dimensionality dependence on the energy equation coming from the diffusion term.
4.2 Mean convergence
As it is clear from (20), the properties of the selection operator $G_{\overline{\mathcal{F}}}[\cdot]$ are crucial for the long time behaviour of the averaged PDE model (13). At the algorithmic level, we can expect that if $\overline{\mathcal{F}}$ attains sharp maxima, few agents will be considered fit for being copied and mutated, leading to fast concentration. On the contrary, if $\overline{\mathcal{F}}$ is almost flat, all the agents will considered almost equally fit, leading to a more dissipative dynamics.
In Genetic Algorithms, this property of the evolution is called selection pressure (Cerf, 1998; Haasdijk and Heinerman, 2018). In our settings, since the weights in $G_{\overline{\mathcal{F}}}[\cdot]$ are proportional to $\exp(\overline{\mathcal{F}})$, the selection pressure can be tuned by introducing an additional parameter $\alpha$ to rescale the fitness as
$\overline{\mathcal{F}}\leftarrow\alpha\overline{\mathcal{F}}\qquad\textup{with}\quad\alpha>0\,.$
Large $\alpha$’s correspond to high selection pressure, while lower ones to a low selection pressure.
In this section, we show that under suitable assumptions on $\overline{\mathcal{F}}$ (in particular, it attains a unique maximizer) and if the selection pressure $\alpha$ is sufficiently high, solutions to the averaged PDE models (13) converge towards the fittest point $h^{\star}$ of problem (17). The analysis technique mirrors the one proposed in (Borghi and Pareschi, 2025) for genetic algorithms. The only differences are that here we are considering a time-continuous dynamics, there is no crossover mechanism, and we are maximizing instead of minimizing.
Assumption 9
The objective function $\overline{\mathcal{F}}:\mathbb{R}^{d_{h}}\to\mathbb{R}$ is continuous and satisfies:
  1. (i)
    (solution uniqueness) there exists a unique global maximizer $h^{\star}$;
  2. (ii)
    (growth conditions) there exists $L_{\overline{\mathcal{F}}},c_{u},c_{l},R_{l}>0$ such that
    $\begin{cases}|\overline{\mathcal{F}}(h)-\overline{\mathcal{F}}(\tilde{h})|\leq L_{\overline{\mathcal{F}}}(1+|h|+|\tilde{h}|)|h-\tilde{h}|&\forall\;h,\tilde{h}\in\mathbb{R}^{d_{h}}\\ \overline{\mathcal{F}}(h^{\star})-\overline{\mathcal{F}}(h)\leq c_{u}(1+|h|^{2})&\forall\;h\in\mathbb{R}^{d_{h}}\\ \overline{\mathcal{F}}(h^{\star})-\overline{\mathcal{F}}(h)\geq c_{u}|h|^{2}&\forall\;h\,:\,|h|>R_{l}\,;\end{cases}$
  3. (iii)
    (inverse continuity) there exists $c_{p},p>0$, $R_{p}>0$ and an upper bound $\overline{\mathcal{F}}_{\infty}>0$ such that
    $\begin{cases}c_{p}|h-h^{\star}|\leq\overline{\mathcal{F}}(h^{\star})-\overline{\mathcal{F}}(h)&\forall\;h\,:\,|h|\leq R_{p}\\ \overline{\mathcal{F}}_{\infty}\leq\overline{\mathcal{F}}(h^{\star})-\overline{\mathcal{F}}(h)&\forall\;h\,:\,|h|>R_{p}\,.\end{cases}$
The first important estimate refer to the effect of the selection operator on the energy of a given distribution $\rho\in\mathcal{P}(\mathbb{R}^{d_{h}})$. Under Assumption 9 (ii) the following estimate holds (Carrillo et al., 2021, Lemma 3.3):
$\mathrm{En}\left(G_{\alpha\overline{\mathcal{F}}}[\rho]\right)\leq b_{1}+b_{2}\mathrm{En}(\rho)\,$, (21)
with
$b_{1}=R_{l}+b_{2}\,,\qquad b_{2}=2\frac{c_{u}}{c_{l}}\left(1+\frac{1}{\alpha c_{l}}\frac{1}{R_{l}^{2}}\right)\,.$
Note that the estimate does not deteriorates as the selection pressure $\alpha$ increases. The next important result we will employ is a quantitative version of the Laplace principle derived in (Fornasier et al., 2024, Proposition 21). Assume $h^{\star}\in\textup{supp}(\rho)$, and let $\overline{\mathcal{F}}_{r}:=\overline{\mathcal{F}}(h^{\star})-\inf_{h\in B(h^{\star},r)}\overline{\mathcal{F}}(h)$, for any $r\in(0,R_{p}]$ and $q>0$ such that $q+\overline{\mathcal{F}}_{r}<\overline{\mathcal{F}}_{\infty}$, we have
$|\textup{m}\left(G_{\alpha\overline{\mathcal{F}}}[\rho]\right)-h^{\star}|\leq c_{p}(q+\overline{\mathcal{F}}_{r})^{1/p}+\frac{\exp(-\alpha q)}{\rho\left(B(h^{\star},r)\right)}\int|h-h^{\star}|\rho(dh)\,.$ (22)
The above estimate relates the mean of the reweighted probability $G_{\alpha\overline{\mathcal{F}}}[\rho]$ with the solution to the optimisation problem (17). Under Assumption 9, it is possible to take the $r>0$ very small such that the first term is smaller with respect to a given tolerance. Then, by taking a sufficiently large selection pressure $\alpha\gg 1$, the second term can also be made arbitrary small. The proof of the following convergence result is based on this intuition, see Appendix C.1 for the details.
Theorem 10
Consider a weak measure solution $\rho\in C([0,t_{\max}],\mathcal{P}_{2}(\mathbb{R}^{d_{h}}))$ to (13), and let the effective fitness $\overline{\mathcal{F}}$ satisfy Assumption 9. If $h^{\star}\in\mathrm{supp}(\rho_{0})$, then, for any tolerance $tol>0$, there exists a selection pressure $\alpha=\alpha(\overline{\mathcal{F}},\rho_{0},t_{\max})>0$ sufficiently large such that
$|\textup{m}(\rho(t))-h^{\star}|^{2}\leq e^{-t}|\textup{m}[\rho_{0}]-h^{\star}|^{2}+(1-e^{-t})\frac{tol}{2}\qquad\textup{for all}\quad t\in[0,t_{\max}]\,.$
Moreover, if
$t_{\max}=t^{\star}\geq\log\left(\frac{2|\textup{m}(\rho_{0})-h^{\star}|^{2}}{tol}\right)$
it holds
$|\textup{m}(\rho(t^{\star}))-h^{\star}|^{2}\leq tol\,.$
Remark 11
In the present work we focus on the convergence scenario associated with a single optimum of the effective fitness, which, intuitively, leads to a unimodal equilibrium distribution around the fittest solution. In more general settings, though, one can expect a richer long-time behaviour, with potentially non-unique, multimodal steady states, and phase-transition-type phenomena as selection pressure varies. The relation discussed in Section 3.2 with descent dynamics with respect to the Hellinger–Kantorovich (Wasserstein–Fisher–Rao) geometry suggests the availability of Lyapunov functionals, and a variational characterisation of steady states. A systematic analysis of these entropy-based mechanisms and possible phase transitions is beyond the scope of the present paper and will be addressed in future work.
5 Numerical experiments
In this section, we numerically investigate the many-agent limit and the emergence of two time scales in the training-evolutionary dynamics. We begin with low-dimensional examples and then validate the proposed framework on a Reinforcement Learning (RL) task. For the illustrative examples, we consider a quadratic bi-level optimisation problem, for which the equilibria can be computed exactly, allowing us to simulate both the limiting dynamics as $N\to\infty$ and the two-time-scale regime $(\varepsilon\to 0)$. We then consider the Himmelblau function, which exhibits a more complex landscape, to study the interplay between parameter training and hyperparameter evolution under Population-Based Training (PBT) updates. Finally, through the deep RL example of the CartPole problem, we illustrate the effect of the two-time-scale structure via effective fitness evaluations and the evolution of the hyperparameter distribution over time. Since sampling parameters from the equilibrium distribution of the parameters is not possible in this setting, we average the fitness over a time window. Assume we need to evaluate the effective fitness at a step $k$ of a neural network $(\theta_{k}^{i},h^{i}_{k})$, this can be approximated as
$\overline{\mathcal{F}}(h_{k})\approx\frac{1}{m}\sum_{s=1}^{m}\mathcal{F}(\theta^{i}_{k-s},h^{i}_{k})\,,\qquad\textup{for a time window}\quad m\geq 1\,$, (23)
where $\theta^{i}_{k-s},s=1,\dots,m$ are the last iterations of the training algorithm (Adam optimizer in our case). Note that, for relatively small time windows $m$, we have $h^{i}_{k}=h^{i}_{k-s}$ since the hyperparameters are updated at a slower rate. Our experiments suggests that larger averaging windows yield more stable and accurate results, supporting the interpretation that the hyperparameter PBT update does not directly optimize the instantaneous fitness, but rather an effective fitness induced by the parameter training equilibrium.
5.1 Illustrative examples
We consider two bi-level optimisation problems in which the fitness function depends only on the parameters, namely $\mathcal{F}=\mathcal{F}(\theta)$ with $\theta\in\mathbb{R}^{2}$. The hyperparameter space is also two-dimensional, with $h=(h_{0},h_{1})\in\mathbb{R}^{2}$. The first hyperparameter, $h_{0}$, will be used to introduce a bias in the loss function $\mathcal{L}=\mathcal{L}(\theta,h_{0})$ used for training. The second hyperparameter, $h_{1}$, determines the noise strength in the SGD training dynamics
$d\theta_{t}=-\nabla_{\theta}\mathcal{L}(\theta,h_{0})\,dt+h_{1}\,dB_{t}\,.$ (24)
Under this choice, the optimal hyperparameters are given by $h_{0}=0$, while $h_{1}$ should decrease over the course of the computation. With these settings, the agents $\{(\theta^{i},h^{i})\}_{i=1}^{N}$ evolve according to Algorithm 1 with $\tau=1$ and mutation noise $\sigma=0.1$. Both parameters and hyperparameters are initially sampled uniformly over [1,1][-1,1]. In the PBT update of the hyperparameters, the agents to be copied are selected proportionally to $\exp(\alpha\mathcal{F}(\theta^{i}))$, with selection pressure $\alpha=100$. This strategy is also known as Gibbs or softmax-based selection. The parameters evolution (24) is discretized via explicit Euler–Maruyama scheme with time step $\Delta t=0.01$.
Quadratic function
We consider a quadratic fitness, similar to the one in (Jaderberg et al., 2017)
$\mathcal{F}(\theta)=1.2-(\theta_{0}^{2}+\theta_{1}^{2})\,,\quad\theta=(\theta_{0},\theta_{1})\in\mathbb{R}^{2},$
which attains its maximum $1.2$ at $\theta=(0,0)$. The loss function corresponds to a bias fitness
$\mathcal{L}(\theta,h_{0}):=-\mathcal{F}(\theta-h_{0})=-1.2+\bigl((\theta_{0}-h_{0})^{2}+(\theta_{1}-h_{0})^{2}\bigr)\,.$
Note that $\mathcal{L}(\cdot,h_{0})$ is also quadratic, and so (24) admits a unique Gaussian invariant measure,
$\mu^{\infty}_{\mathcal{L}}(\theta\mid h)=\frac{e^{-2\mathcal{L}(\theta,h_{0})/{h_{1}^{2}}}}{\int e^{-{2}\mathcal{L}(\theta,h_{0})/{h_{1}^{2}}}d\theta}=\mathcal{N}\!\left(\begin{pmatrix}h_{0}\\ h_{0}\end{pmatrix},\frac{h_{1}^{2}}{4}I_{2}\right).$ (25)
(a) Fitness
(b) Loss
Figure 2: Quadratic objective fitness and loss. Median and $10\%-90\%$ quantiles of the population fitness $\mathcal{F}(\theta)$ and loss $\mathcal{L}(\theta,h)$ over $100$ iterations of Algorithm 1. PBT updates, where agents are resampled according to their fitness values, occur every 50 training steps and lead to a discontinuity in the fitness evolution. Population size is $N=100$.
(a) Evolution of $h_{0}$ marginal distribution
(b) Evolution of $h_{1}$ marginal distribution
Figure 3: Many-agent limit $N\to\infty$. Evolution of the hyperparameter distributions for different population sizes $N=10^{3},10^{4},10^{5}$ and quadratic problem. Agents are initially sampled uniformly from [1,1][-1,1] and evolved according with Algorithm 1. The total number of iterations is $500$, with selection jumps at every 50 training steps. The curves are constructed by connecting bin centres (markers) with straight lines and the number of bins is automatically determined.
(a) Evolution of $h_{0}$ marginal distribution
(b) Evolution of $h_{1}$ marginal distribution
Figure 4: Fast training limit $\varepsilon\to 0$. Evolution of the hyperparameter distributions for different fast-training regimes. The limit $\varepsilon\to 0$ in the PDE model (6) corresponds in Algorithm 1 to increasing the number of internal SGD steps performed between successive PBT hyperparameter updates. We consider 20, 50, and 100 inner training steps. The case $\varepsilon=0$ corresponds to directly sampling from the equilibrium distribution (25) and evolving the agents according to Algorithm 2, which simulates the averaged PDE model (13). The population size is $N=10^{5}$, and the curves are obtained by connecting bin centers with straight lines using 50 bins.
Figure 2 shows how the population fitness and loss evolve during training and across PBT updates with $N=100$ agents. Immediately after each PBT update, the median fitness exhibits a sharp jump toward the optimum, as the top-performing agents are copied. The fitness then gradually decreases until the next update. This occurs because the SGD/Langevin dynamics optimise the biased loss function. When exploration noise is added to the hyperparameters $h$, the Langevin dynamics drive the parameters $\theta$ toward these mutated, suboptimal targets, temporarily worsening the true fitness, while the biased loss continues to decrease.
Figure 3 illustrates the many-agent limit of the PBT dynamics for $N=10^{3},10^{4},10^{5}$. As predicted by Theorem 4, the evolution becomes less noisy as $N$ increases, suggesting convergence toward the deterministic kinetic model (6). Since the mutation strength does not decrease during the computation, the large-time distribution of the hyperparameters is approximately Gaussian and centered around the optimal values $h_{0}=0$ and $h_{1}=0$. This is consistent with the convergence analysis of Theorem 10.
Figure 4 presents the two-time-scale limit in which the SGD training dynamics is rescaled by a parameter $\varepsilon>0$. At the algorithmic level, this corresponds to increasing the number of inner training steps performed between successive evolutionary hyperparameter updates. We consider 20, 50, and 100 training steps per PBT update, corresponding to the rescaled PDE evolution (6) in the regime $\varepsilon\to 0$. The limiting case $\varepsilon=0$, corresponding to the averaged PDE (13), is simulated using Algorithm 2, in which the iterative inner training loops are replaced by direct sampling from the invariant measure (25).
Simulating the averaged equation significantly reduces the computational cost. Indeed, the dominant cost of PBT, the inner-loop training, is replaced by a single sampling step from a Gibbs distribution. Moreover, we observe that, as more training steps are performed between consecutive PBT updates, the hyperparameter distribution more closely matches the one produced by the reduced algorithm. Also, the two-time-scale separation seems not to be uniform in time as the difference between the distributions becomes less evident after 4 PBT updates, see Figure 4 (right).
Algorithm 2 Training via averaged PDE (13)
Initialize $\in\mathbb{N} NN$s with hyperparameters $h^{i}\in H$, $i=1,\dots,N$
Set $\tau\in(0,1]$$\sigma>0$, $t_{\max}>0$
$t\leftarrow 0$
while $t<t_{\max}$ do
  for $i=1,\dots,N$ do
   $\overline{F}^{i}\leftarrow\mathcal{F}(\theta^{i},h^{i})$ with $\theta^{i}\sim\mu^{\infty}(\theta|h^{i})$ $\triangleright$ effective fitness via sampling
   OR:
   $\overline{F}^{i}\leftarrow\overline{\mathcal{F}}(h^{i})$ approximated via (23) $\triangleright$ effective fitness via time averaging
  end for
  for $i=1,\dots,N$ do $\triangleright$ population interaction
   with probability $\tau$:
     select a NN $(\theta^{j},h^{j})$ proportionally to $\exp(\overline{F}^{j})$
     sample random mutation vector $\xi$
     $\tilde{\theta}^{i}\leftarrow\theta^{j}$ $\triangleright$ copy parameter
     $\tilde{h}^{i}\leftarrow h^{j}+\sigma\xi$ $\triangleright$ copy and perturb hyperparameter
  end for
  $\theta^{i}\leftarrow\tilde{\theta}^{i},h^{i}\leftarrow\tilde{h}^{i}$ for all $i=1,\dots,N$
  $t\leftarrow t+\tau$
end while
Himmelblau function
To study parameter distributions in a non-convex setting, we consider the two-dimensional Himmelblau function,
$\mathcal{F}(\theta)=(\theta_{0}^{2}+\theta_{1}-11)^{2}+(\theta_{0}+\theta_{1}^{2}-7)^{2},\quad\theta=(\theta_{0},\theta_{1})\in\mathbb{R}^{2},$ (26)
with, as before, a biased loss for $h_{0}\in\mathbb{R}$
$\mathcal{L}(\theta,h_{0})=((\theta_{0}-h_{0})^{2}+\theta_{1}-11)^{2}+(\theta_{0}+(\theta_{1}-h_{0})^{2}-7)^{2},$ (27)
The Himmelblau function has four distinct global minima, each attaining $\mathcal{F}=0$ (see Figure 5(a)). Since our framework assumes maximisation of the fitness we will consider -$\mathcal{F}$ when evaluating the agents’ fitness.
Figure 5(b) illustrates the underlying mechanism of Algorithm 1: agents first approach nearby minima via SGD dynamics, after which copying high-performing agents can relocate particles into different basins. Subsequent SGD steps refine the parameters within these basins, ultimately leading to concentration around the a single optimum located at $(3,2)$.
(a) Himmelblau function
(b) Evolution of the distribution of $\theta$
Figure 5: Himmelblau test, parameters’ evolution. The population size is $N=10^{5}$, and the parameters $\theta$ are initially sampled uniformly from the box $0.5,0.5]^{2}$. The density is reconstructed using a 40-bin histogram in each dimension. The second row in (b) shows the trajectories of 5 randomly selected agents. The solid orange curves correspond to the SGD training dynamics, while the red dashed lines indicate the jumps induced by the PBT update after an inner training loop.
5.2 Deep Reinforcement Learning
In this test, we show how the two-time-scale modelling framework extends beyond the simplified settings considered in our analysis. We consider a population of neural networks that learn collectively to control a Cart-Pole system through the PBT algorithm. Training is carried out using the Adam optimiser, while the selection-mutation updates are performed via rank-based selection, in which the worst-performing neural networks are replaced with copies of the best-performing ones, see Section 2.4 for more details. We investigate the emergence of a two-time-scale separation by considering different numbers of inner training steps and by approximating the effective fitness through time averaging, as in (23).
Settings.
Controlling a Cart-Pole system is one of the classic benchmark problems in control theory and reinforcement learning (Barto et al., 1983; Sutton and Barto, 1998). In this environment, a pole is attached to a cart via an unactuated joint, and the cart can move along a frictionless track. The pole starts in an upright position, and the objective is to keep it balanced for as long as possible by applying forces to the cart either to the left or to the right.
The state at time steps $t=0,1,2,\dots$ consists of four components:
$s_{t}=(x_{t},\dot{x}_{t},\phi_{t},\dot{\phi}_{t}),$
which are, respectively, the cart position, cart velocity, pole angle, and pole angular velocity. The action space contains two discrete actions, $a_{t}$: left ($a_{t}=0$) and right ($a_{t}=1$). The cart position $x$ lies within the interval (4.8,4.8)(-4.8,4.8), and an episode terminates early if the cart leaves the range (2.4,2.4)(-2.4,2.4). Similarly, the pole angle $\phi$ is observable within (0.418,0.418)(-0.418,0.418) radians, and the episode ends if $\phi\notin(-0.2095,0.2095)$.
By default, the agent receives a reward $r_{t}=1$ at every time step, including the termination step, reflecting the objective of maximizing the duration for which the pole remains upright. An episode concludes either prematurely, when a failure condition is met (the pole falls or the cart goes out of bounds), or successfully, when the cumulative reward reaches a predefined maximum, which can be adjusted.
To investigate the two-time-scale dynamics of neural network training, we implement a Deep Q-Network (DQN) (Mnih et al., 2015). The network architecture consists of an input layer with four state nodes, two hidden layers with 64 units each, and an output layer with two action nodes. Here, we use Adam (Kingma and Ba, 2014) for training, and PBT is employed to optimize three hyperparameters: the learning rate of the Adam optimiser, the greedy policy decay rate, and the experience replay batch size. To select the best agents to be copied, we use a rank-based strategy, also known as truncation selection (see Section 2.4). Agents are assigned a rank according to their fitness, and the top $20\%$ of the population is copied to replace the bottom $20\%$. Further algorithmic and hyperparameter details are provided in Appendix D.
Results.
Our modelling analysis suggested that, in the presence of a two-time-scale separation, the slow hyperparameter dynamics attempts to optimise a fitness that is averaged with respect to the parameter distribution. In this experiment, the fast time scale corresponds to the environment interactions (RL training steps), while the slow time scale corresponds to the evolutionary hyperparameter updates. Since sampling from equilibrium is not possible, we instead perform a time averaging of the fitness, which is here the episodic rewards.
We define a generation as a single PBT update interval consisting of $300,500$, or $1000$ training steps. The maximum episodic reward is capped at $100$, so an agent completes at least $3,5$, and $10$ full episodes before evaluation, respectively. During the slow-scale update, an agent’s fitness score is calculated as the moving average of its last $m$ episodic rewards. Figure 6 illustrates that increasing the moving-average window size $m$ effectively smooths the fast-scale performance noise.
Figure 7 shows that, after $30$ PBT updates, high-performing hyperparameter configurations concentrate in a localised region of the search space, supporting mean convergence toward the fittest solutions (Theorem 10). The influence of the mutation noise $\sigma$ is also evident: larger values of $\sigma$ lead to a more dispersed distribution of hyperparameters across the population.
(a) 300 training steps
(b) 500 training steps
(c) 1000 training steps
Figure 6: Cart–Pole control: two-time-scale dynamics. Mean episodic reward of top five agents (solid line with markers) and entire population (dashed line) across generations. We solve the RL problem with $N=500$ NN agents via Population-Based Training evolution. The parameter averaging window size indicates the number of recent episodes ($m$) averaged to compute the agent’s fitness score. Larger window sizes yield more stable evaluations, reducing fast-scale noise and leading to faster convergence.
(a) Low mutation noise, $\sigma=0.05$
(b) High mutation noise, $\sigma=0.1$
Figure 7: Cart–Pole control: hyperparameter distribution. For two different values of the mutation noise $\sigma$, we show the distribution of hyperparameters across $N=500$ agents, using $1000$ training steps and an averaging window of size $m=2$. The maximum episodic reward is $500$. The marginal distributions in each panel are estimated via Gaussian kernel density estimation.
6 Conclusion and outlook
The results of this work suggest a different way to understand, and possibly design, population-based neural training. By placing heuristic population methods within a continuous-time large-population framework, we move from viewing training as a collection of algorithmic rules toward a description in which learning is shaped by collective statistical dynamics. In the strong separation regime, the joint evolution reduces to a closed selection–mutation equation for the hyperparameter density $\rho(t,h)$ driven by the Gibbs-averaged fitness $\overline{\mathcal{F}}(h)$. In this setting, population-level quantities, rather than individual optimisation trajectories, become the natural objects of analysis, comparison, and control.
A key consequence of this perspective is the emergence of an effective fitness as a macroscopic, statistically averaged quantity. In the asymptotic regime considered here, performance is no longer tied to the outcome of a single training run, but to properties of an equilibrium distribution induced by noisy learning dynamics. This suggests the possibility of surrogate training strategies in which explicit optimisation of individual networks is partially replaced by analytically or numerically accessible population-level criteria. While such ideas remain to be explored beyond the idealised regimes considered here, they point toward a principled route to reducing computational cost, improving stability, and increasing interpretability in complex training pipelines.
The framework developed here makes it possible to study population-based learning as a dynamical system, using questions familiar from kinetic theory and statistical physics, including stability, concentration, phase transitions, and the role of noise and interaction in shaping macroscopic behaviour. It also provides a natural basis for the systematic design of meta-learning and hyperparameter optimisation protocols, including optimal control formulations on the reduced dynamics.
Extending this approach to heterogeneous populations, structured interactions, crossover or model-merging mechanisms, finite-population effects, and non-equilibrium training scenarios is a natural direction for future research, with potential implications for distributed, federated, and large-scale ensemble learning.
The replicator–mutator structure of the averaged dynamics suggests that the effectiveness of PBT is not entirely heuristic: the convergence guarantee to the optimal hyperparameters configuration provides a principled basis for the selection pressure and mutation noise design choices made in practice. Developing these ideas into a quantitative theory may help make population-based training a more systematic and interpretable component of modern machine learning.
Aknowledgments
The work of Lorenzo Pareschi and Giacomo Borghi was supported by the Royal Society under the Wolfson Fellowship “Uncertainty quantification, data-driven simulations and learning of multiscale complex systems governed by PDEs”. Lorenzo Pareschi also acknowledges the partial support of the FIS2023-01334 Advanced Grant “Tackling complexity: advanced numerical approaches for multiscale systems with uncertainties” (ADAMUS).
Appendix A Notation
[…]
With $\mathcal{P}(\mathbb{R}^{d})$ we indicate the set of Borel probability measures over $\mathbb{R}^{d}$, and with $\mathcal{P}_{p}(\mathbb{R}^{d})$ those with bounded $p$-th moment, $M_{p}(f):=\int|x|^{p}f(dx)<\infty$. We will sometime use the shorter notation
$\langle\phi,f\rangle:=\int\phi(x)f(dx)$
to indicate integrals for a measure $f$ and a measurable test function $\phi$. We equip $\mathcal{P}_{p}(\mathbb{R}^{d})$ with the $L^{p}$-Wasserstein distance (Villani, 2009)
$\mathbb{W}_{p}(f,g):=\left(\mathbb{E}_{X\sim f,Y\sim g}|X-Y|^{p}\right)^{1/p}\,.$
Appendix B Proof of Theorem 4: propagation of chaos
The proof follows the technique proposed in (Borghi, 2026) to study the many agent limit of genetic algorithm. Here, we need to additionally consider the SGD training in the parameters space.
For notational simplicity, we will use the stack parameters $\theta\in\mathbb{R}^{d_{\theta}}$, and hyperparameters $h\in\mathbb{R}^{d_{h}}$ in single vector $x:=(\theta,h)\in\mathbb{R}^{d}$ with $d:=d_{\theta}+d_{h}$. Note that Algorithm 1 divides the evolution of the $N$ neural networks agents $X^{i}=(\theta^{i},h^{i})$ into intervals of length $\tau\in(0,1]$. Inside the intervals, agents evolve according to the gradient-based training dynamics, while at $t=t_{n}$, they are subject to a genetic PBT update. Let $t_{n}=n\tau$, and consider the function $\bm{b}(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d}$ and matrices $\bm{\Sigma},\bm{\sigma}\in\mathbb{R}^{d\times d}$
$\bm{b}(x):=\begin{pmatrix}-\nabla_{\theta}\mathcal{L}(\theta,h)\\ \bm{0}_{d_{h}}\end{pmatrix}\,,\quad\bm{\Sigma}:=2\beta^{-1}\mathrm{diag}\begin{pmatrix}\bm{1}_{d_{\theta}}\\ \bm{0}_{d_{h}}\end{pmatrix}\quad\textup{and}\quad\bm{\sigma}:=\sigma\mathrm{diag}\begin{pmatrix}\bm{0}_{d_{\theta}}\\ \bm{1}_{d_{h}}\end{pmatrix}\,$.
For a given probability measure $f\in\mathcal{P}(\mathbb{R}^{d})$, we write the selection operator for the fitness as
$G_{\mathcal{F}}[f](dx):=\frac{e^{\mathcal{F}(x)}f(dx)}{\langle e^{\mathcal{F}},f\rangle}\,.$
The evolution for $(X_{t}^{i})_{t\geq 0}$ described by Algorithm 1, can then be re-written as following. Starting from some initial data $X_{0}^{i}\sim f_{0}$, for each iteration $n=0,1,\dots$, in the open interval $(t_{n},t_{n+1})$ the agent is defined as the solution to the Cauchy problem
$\begin{dcases}dX_{t}^{i}=\bm{b}(X_{t}^{i})dt+\bm{\Sigma}^{1/2}dW^{i}_{t}\qquad t\in(t_{n},t_{n+1})\\ \lim_{t\to t_{n}^{+}}X_{t}^{i}=X_{t_{n}}^{i}\end{dcases}\qquad i=1,\dots,N\,$, (28)
while at the discrete times $t_{n+1}$, agents undergo a resampling-mutation procedure. Recall that an agent $j$ is selected to be copied with a probability $w^{j}=\exp(\mathcal{F}(X^{j}))/\sum_{\ell}\exp(\mathcal{F}(X^{\ell}))$. One can implement it as follows: given a set of weights $\mathbf{w}\in\Delta_{N}$, define the index maps $\mathbf{j}(\mathbf{w}.\cdot):[0,N)\to\{1,\dots,N\}$ as
$\mathbf{j}(\mathbf{w},\alpha):=\min\left\{j\in\{1,\dots,N\}\,\middle|\,\alpha<N\sum_{\ell=1}^{j}w^{\ell}\right\}\,.$
Note that if $\alpha\sim[0,N)$, then $\mathbb{P}(\mathbf{j}(\mathbf{w},\alpha)=j)=w^{j}$. Consider i.i.d. random variables $\tau^{i}_{n}\sim\textup{Bern}(\tau)$, $\alpha^{i}\sim\textup{Unif}[0,N)$, $\xi_{n}^{i}\sim\mathcal{N}(\bm{0}_{d},I_{d})$, and let $\mathbf{w}_{n}:=\exp(\mathcal{F}(X_{t_{n+1}^{-}}^{j}))/\sum_{\ell}\exp(\mathcal{F}(X_{t_{n+1}^{-}}^{j})$. The genetic update can then be written as
$X^{i}_{t_{n+1}}=(1-\tau_{n}^{i})X^{i}_{t_{n+1}^{-}}+\tau_{n}^{i}\left(X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i}_{n})}_{t_{n+1}^{-}}+\bm{\sigma}\xi^{i}_{n}\right)\,.$ (29)
Recall the limiting kinetic PDE model (6) is given by
[…]
where the training and evolutionary operator are weakly defined, in our notation, as
$\langle\phi,T_{\mathcal{L}}[f]\rangle$ $:=\int\left(\phi(x)\cdot\bm{b}(x)+\frac{1}{2}\nabla\cdot(\bm{\Sigma}\nabla\phi)\right)f(dx)$
$\langle\phi,E_{\mathcal{F}}[f]\rangle$ $:=\int\phi(\tilde{x}+\bm{\sigma}\xi)\,G_{\mathcal{F}}[f](d\tilde{x})\,\mu^{\xi}(d\xi)-\int\phi(x)f(dx)\,.$
Definition 12 (Weak measure solution)
Let $f_{0}\in\mathcal{P}_{q}(\mathbb{R}^{d})$ be an initial datum. We say that $f\in C([0,T],\mathcal{P}_{q}(\mathbb{R}^{d}))$ is a weak measure solution to the Cauchy problem (30) if for any $\phi\in C_{0}^{\infty}(\mathbb{R}^{d})$ it holds
$\frac{d}{dt}\langle\phi,f(t)\rangle=\langle\phi,T_{\mathcal{L}}[f(t)]+\langle\phi,E_{\mathcal{F}}[f(t)]\rangle$
for almost every $t\in[0,t_{\max}]$, and $\lim_{t\to 0^{+}}\langle\phi,f(t)\rangle=\langle\phi,f_{0}\rangle$.
B.1 Preliminaries
We introduce an equivalent formulation of the BL norm (7), as an optimal transport problem. Let $f,g\in\mathcal{P}(\mathbb{R}^{d})$. Thanks to the Kantorovich–Rubinstein duality (Villani, 2009, Theorem 1.14), it holds
$\|f-g\|_{\mathrm{BL}}=\inf_{X\sim,Y\sim g}\mathbb{E}\left[|X-Y|\wedge 1\right]\,.$ (31)
That is, the Wasserstein-1 distance with cost $D^{\wedge}(x,y):=|X-Y|\wedge 1$ admits a dual representation which corresponds to the BL norm, see (Hairer and Mattingly, 2008; Piccoli and Rossi, 2016; Borghi, 2026) for more details. An interesting consequence is that, since $D^{\wedge}(x,y)\leq|x-y|$, $\|\cdot\|_{\mathrm{BL}}$ is upper bounded by the usual Wasserstein-1 distance $\mathbb{W}_{1}$ with cost |xy||x-y|,
$\|f-g\|_{\mathrm{BL}}\leq\mathbb{W}_{1}(f,g)\,.$
We collect some auxiliary results that will be useful later.
Lemma 13 (Concentration rate)
Let $f\in\mathcal{P}_{q}(\mathbb{R}^{d})$ with $q>1$ as in Assumption 3, and $f^{N}=(1/N)\sum_{i=1}^{N}\delta_{X^{i}}$, with $X^{i}\sim f$ i.i.d.. There exists a constant $C=C(d,q)>0$ such that for all $N\geq 1$:
$\|f-f^{N}\|_{\mathrm{BL}}\leq CM_{q}^{1/q}(f)\,\epsilon(N)$
with $\epsilon(N)\to 0$ as $N\to\infty$.
Proof The inequality holds for $\mathbb{W}_{1}(f,f^{N})$ as a special case of (Fournier and Guillin, 2015, Theorem 1). Then, we use that $\|f-f^{N}\|_{\mathrm{BL}}\leq\mathbb{W}_{1}(f,f^{N})$. The explicit form of the rate $\epsilon(N)$ depends on $q,d$, and can be found in (Fournier and Guillin, 2015).
Lemma 14 (Selection stability (Borghi, 2026, Lemma 2.4))
Let $\mathcal{F}$ be bounded, globally Lipschitz and strictly positive. There exists a positive constant $C=C(\mathcal{F})$ such that for any $f,g\in\mathcal{P}(\mathbb{R}^{d})$ it holds
$\|G_{\mathcal{F}}[f]-G_{\mathcal{F}}[g]\|_{{\mathrm{BL}}}\leq C_{\mathcal{F}}\|f-g\|_{\mathrm{BL}}\,.$
B.2 Limit $N\to\infty$
We aim to prove propagation of chaos via coupling methods, where the system of interacting agents $\mathbf{X}_{t}:=(X_{t}^{1},\dots,X_{t}^{N})$ is coupled with a system of non-linear, non-interacting agents $\overline{\mathbf{X}}^{i}_{t}:=(\overline{X}_{t}^{1},\dots,\overline{X}_{t}^{N})$. Then, we aim to estimate their distance by using the optimal transport formulation of the BL norm (31).
Non-linear particle system.
Initial data is the same: $\mathbf{X}_{0}=\overline{\mathbf{X}}_{0}$. Then, as before, the non-linear system evolves in the time interval $(t_{n},t_{n+1})$ as
$\begin{dcases}d\overline{X}_{t}^{i}=\bm{b}(\overline{X}_{t}^{i})dt+\bm{\Sigma}^{1/2}dW^{i}_{t}\qquad t\in(t_{n},t_{n+1})\\ \lim_{t\to t_{n}^{+}}\overline{X}_{t}^{i}=\overline{X}_{t_{n}}^{i}\end{dcases}\quad i=1,\dots,N\,$,
with the same Brownian paths $(W_{t}^{i})_{t\geq 0}$ of the system $\mathbf{X}_{t}$.
Let $f^{\tau}(t):=\textup{Law}(\overline{X}_{t}^{i})$. To define the genetic PBT update, we consider a measurable function $X^{*}_{n}:(\mathbb{R}^{d})^{N}\times\textup{int}(\Delta_{N})\times[0,N)\to\mathbb{R}^{d}$ with the following property: if $\alpha\sim\textup{Unif}[0,N)$, then
$\textup{Law}\left(X^{*}\left(\mathbf{w}_{n},\mathbf{X}_{t_{n+1}^{-}},\alpha\right)\right)=G_{\mathcal{F}}\left[f(t_{n+1}^{-})\right]\,$,
and
$\mathbb{E}_{\alpha}\left[\left|X_{n}^{*}\left(\mathbf{w}_{n},\mathbf{X}_{t_{n+1}^{-}},\alpha\right)-X^{\mathbf{j}(\mathbf{w}_{m},\alpha_{n}^{i})}\right|\wedge 1\right]=\left\|G_{\mathcal{F}}\left[f^{\tau}(t_{n+1}^{-})\right]-G_{\mathcal{F}}\left[f^{N,\tau}(t_{n+1}^{-})\right]\right\|_{{\mathrm{BL}}}\,.$
Existence of such map is ensured by (Borghi, 2026, Lemma 3.1) under Assumption 3.
The genetic update for each non-linear particle is then given by
$\overline{X}^{i}_{t_{n+1}}=(1-\tau_{n}^{i})\overline{X}^{i}_{t_{n+1}^{-}}+\tau_{n}^{i}\left(X_{n}^{*}(\mathbf{X}_{t},\mathbf{w}_{n},\alpha^{i}_{n})+\bm{\sigma}\xi^{i}_{n}\right)\,$, (32)
where we used the same random variables $\alpha^{i}_{n},\tau_{n}^{i},\xi_{n}^{i}$ of (32). The evolution of $\overline{X}_{t}^{i}$ is of McKean–Vlasov type since it depends on the law $f^{\tau}(t)$ of the agent itself. Also, the agents are interdependent on one another since the auxiliary random variables $\alpha^{i}_{n},\tau_{n}^{i},\xi_{n}^{i}$ are all sampled independently.
Lemma 15 (Moments bound)
Assume $f_{0}\in\mathcal{P}_{q}(\mathbb{R}^{d})$, and Assumption 3 holds for $\mathcal{L},\mathcal{F}$. Then $f^{\tau}(t)=\textup{Law}(\overline{X}_{t}^{i})$ satisfies
$M_{q}^{1/q}(f^{\tau}(t))\leq C\left(1+M_{q}^{1/q}(f_{0})\right)e^{C^{\prime}t}\qquad\textup{for all}\quad t\in[0,t_{\max}]\,$,
for some positive constants $C,C^{\prime}$ depending on $\mathcal{L},\mathcal{F},\beta,\sigma,q$.
Proof By Assumption 3, $\mathcal{L}$ is globally Lipschitz and so the drift coefficient in the SDE (28) is Lipschitz and admits linear growth. By standards results in SDE theory, see for instance (Mao, 2011, Theorem 4.1), it holds for $q\geq 2$
$\mathbb{E}|\overline{X}_{t}^{i}|^{q}\leq C^{\prime}(1+\mathbb{E}|\overline{X}_{t_{n}}^{i}|^{q})e^{C(t-t_{n})}\quad\textup{for all}\quad t\in[t_{n},t_{n+1})$
with $C(\mathcal{L},\beta,q), C=C$$^{\prime}=C^{\prime}(q)>0$. This leads to $M_{q}(f^{\tau}(t^{-}_{n+1}))\leq C^{\prime}(1+M_{1}(f^{\tau}(t_{n})))e^{C\tau}$. Since by Assumption 3 the fitness function $\mathcal{F}$ satisfies $0<\inf\mathcal{F}\leq\mathcal{F}(x)\leq\sup\mathcal{F}<\infty$ for all $x\in\mathbb{R}^{d}$, we have for some constant $\tilde{C}$ depending on $\sigma,q$
$\int|x|^{q}f^{\tau}(t_{n+1},dx)$ $=(1-\tau)M_{q}(f^{\tau}(t_{n+1}^{-}))+\tau\int|x+\bm{\sigma}\xi|^{q}G_{\mathcal{F}}[f^{\tau}(t_{n+1}^{-})](dx)$
$\leq(1-\tau)M_{q}(f^{\tau}(t_{n+1}^{-}))+\tau\tilde{C}\left(1+\int\frac{e^{\mathcal{F}(x)}|x|^{q}}{\langle e^{\mathcal{F}},f^{\tau}(t_{n+1}^{-})\rangle}f^{\tau}(t_{n+1}^{-},dx)\right)$
$\leq(1-\tau)M_{q}(f^{\tau}(t_{n+1}^{-}))+\tau\tilde{C}\left(1+\exp(\sup{\mathcal{F}}-\inf{\mathcal{F}})M_{q}(f^{\tau}(t_{n+1}^{-}))\right)\,.$
Therefore, we obtain the estimate for some constants $C,C^{\prime}$ depending on $\mathcal{L},\mathcal{F},\sigma,\beta,q>0$
$M_{q}(f^{\tau}(t_{n+1}))\leq(1+C\tau)\max\left\{M_{1}(f^{\tau}(t_{n+1}^{-})),1\right\}\leq e^{C^{\prime}\tau}\max\{M_{q}(f^{\tau}(t_{n})),1\}$
where we further used that $1+s\leq e^{s}$. The desired estimate can be obtain by iterating the argument until $t_{n}=t_{\max}$.
Grönwall-type argument.
We now consider the empirical measure $\overline{f}^{N,\tau}(t)$ associated with the non linear system $\overline{\textbf{X}}_{t}$, and track its distance from $f^{N,\tau}(t)$. Recall that since $\overline{f}^{N,\tau}(t)$ is made of independent i.i.d particles with law $f^{\tau}(t)$, and so $\|\overline{f}^{N\tau}(t)-f^{\tau}(t)\|_{\mathrm{BL}}\lesssim\epsilon(N)$ by Lemma 13.
Since the Brownian paths are shared among the two systems, we have for $t\in[t_{n},t_{n+1})$
$d|X_{t}^{i}-\overline{X}_{t}^{i}|^{2}$ $=2(X_{t}^{i}-\overline{X}_{t}^{i})\cdot(\bm{b}(X_{t}^{i})-\bm{b}(\overline{X}_{t}^{i}))dt$
$\leq 2\|\mathcal{L}\|_{\textup{Lip}}|X_{t}^{i}-\overline{X}_{t}^{i}|^{2}dt\,.$
By Grönwall’s inequality, we obtain $|X_{t}^{i}-\overline{X}_{t}^{i}|^{2}\leq\exp(2\|\mathcal{L}\|_{\textup{Lip}}(t-t_{n}))|X_{t_{n}}^{i}-\overline{X}_{t}^{i}|^{2}$ from which also follows
$\left(|X^{i}_{t}-\overline{X}^{i}_{t}|\wedge 1\right)\leq e^{\|\mathcal{L}\|_{\textup{Lip}}(t-t_{n})}\left(|X^{i}_{t_{n}}-\overline{X}^{i}_{t_{n}}|\wedge 1\right)\qquad\textup{for}\quad t\in[t_{n},t_{n+1})\,.$ (33)
Next, by definition of $X_{t_{n+1}}^{i},\overline{X}_{t_{n+1}}^{i}$, and in particularly their optimal coupling with respect to $D^{\wedge}(x,y)=|x-y|\wedge 1$ we have
$\mathbb{E}\left(|X_{t_{n+1}}^{i}-\overline{X}_{t_{n+1}}^{i}|\wedge 1\right)$ $=(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)+\tau\mathbb{E}\left(|X_{n}^{*}\left(\mathbf{w}_{n},\mathbf{X}_{t_{n+1}^{-}},\alpha\right)-X^{\mathbf{j}(\mathbf{w}_{m},\alpha_{n}^{i})}|\wedge 1\right)$
$=(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)+\tau\left\|G_{\mathcal{F}}\left[f^{N,\tau}(t_{n+1}^{-})\right]-G_{\mathcal{F}}\left[f^{\tau}(t_{n+1}^{-})\right]\right\|_{\mathrm{BL}}\,$
$\leq(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)+\tau C_{\mathcal{F}}\left\|f^{N,\tau}(t_{n+1}^{-})-f^{\tau}(t_{n+1}^{-})\right\|_{\mathrm{BL}}\,$
by Lemma 14. To bound the second term, we apply triangular inequality together with Lemmas 13 and 15, and obtain
$\|f^{N,\tau}(t_{n+1}^{-})-f^{\tau}(t_{n+1}^{-})\|_{\mathrm{BL}}$ $\leq\|f^{N,\tau}(t_{n+1}^{-})-\overline{f}^{N,\tau}(t_{n+1}^{-})\|_{\mathrm{BL}}+\|\overline{f}^{N,\tau}(t_{n+1}^{-})-f^{\tau}(t_{n+1}^{-})\|_{\mathrm{BL}}$
$\leq\frac{1}{N}\sum_{i=1}^{N}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)+C(1+M^{1/q}_{q}(f_{0}))e^{C^{\prime}t_{n+1}}\epsilon(N)\,.$
Summing for all $i=1,\dots,N$ and applying (33), we get the estimate
$E_{n+1}:$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left(|X_{t_{n+1}}^{i}-\overline{X}_{t_{n+1}}^{i}|\wedge 1\right)$
$\leq(1-C\tau)e^{\|\mathcal{L}\|_{\textup{Lip}}\tau}E_{n}+\tau C(1+M^{1/q}_{q}(f_{0}))\epsilon(N)e^{C^{\prime}t_{n+1}}$
$\leq e^{C\tau}E_{n}+\tau C(1+M^{1/q}_{q}(f_{0}))\epsilon(N)e^{C^{\prime}t_{n+1}}\,.$
By iterating for all $n$, until $t_{n}=t_{\max}$, one obtains since $E_{0}=0$
$E_{n}$ $\leq e^{Ct_{n}}E_{0}+\tau C(1+M^{1/q}_{q}(f_{0}))\sum_{k=0}^{n-1}e^{C^{\prime}t_{n}}(1+C\tau)^{n}$
$\leq Ct_{n}(1+M^{1/q}_{q}(f_{0}))e^{C^{\prime}t_{n}}\epsilon(N)\,.$
where the constants depend on $\mathcal{L},\mathcal{F},\beta,\sigma$, and $q$. By the optimal transport formulation of the BL norm (31), we finally get
$\sup_{t\in[0,t_{\max}]}\|f^{N,\tau}(t)-f^{\tau}(t)\|_{{\mathrm{BL}}}\;\;\lesssim\;\;\epsilon(N)\,.$ (34)
where the hidden constant depends also on $t_{\max}$, but not on $\tau$.
B.3 Limit $\tau\to 0$
We study the limit towards the time-continuous model by first showing that $(f^{\tau})_{\tau>0}$ admits a converging subsequence, and then by identifying the limit with a weak measure solution to (30). In the following, $(\overline{X}_{t}^{\tau})_{t\in[0,t_{\max}]}$ is an an arbitrary non-linear particle (defined in the previous section), where the apex $\tau$ stresses the dependence on the time-discretisation parameter. We have $\textup{Law}(\overline{X}_{t}^{\tau})=f^{\tau}(t)$.
Existence of the limit.
Note that $\overline{X}^{\tau},\tau>0,$ are not continuous processes due to the jump at times $t_{n}=n\tau$, but still they are right-continuous and the left limit always exists. Therefore, we consider the topology given by the Skorokhod space of cádlág functions, and show that the sequence is compact and tight. This approach is standard, and we refer to the surveys (Chaintron and Diez, 2022a, b) for a reminder on Skorokhod topology in the context chaos propagation results. We outline the main arguments below.
First, we note that the moments bound derived in Lemma 16 is independent of $\tau$. Therefore, the $q$-th moments of $f^{\tau}(t)$ are uniformly bounded for all $t\in[0,t_{\max}]$ and $\tau>0$. Then, while $f^{\tau}$ is not continuous due to the jumps, the jump size is determined by the parameter $\tau$. Also, we have for every
$\mathbb{E}|\overline{X}^{\tau}_{t}-\overline{X}^{\tau}_{s}|\lesssim|s-t|+\sqrt{|t-s|}\qquad\textup{for any}\quad[s,t]\subset[t_{n},t_{n+1})\,$,
thanks to the globally Lipschitz continuity of the drift $\bm{b}$ and the constant diffusion coefficients $\bm{\Sigma}^{1/2}$. Therefore any sequence $(\tau_{k})_{k\in\mathbb{N}}$ such that $\tau_{k}\to 0$ leads to a sequence $(\overline{X}^{\tau_{k}}_{t})_{t\in[0,t_{\max}]}$, $k\in\mathbb{N}$, which satisfies Assumption 12, 13, and 14 in (Chaintron and Diez, 2022b). By (Chaintron and Diez, 2022b, Theorem C.5), then, we have existence of a limit $f\in C([0,t_{\max}],\mathcal{P}(\mathbb{R}^{d}))$.
Limit is a solution to the PDE.
Fix any $\tau>0$ and take $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. It holds
$\langle\phi,f^{\tau}(t_{n+1})-f^{\tau}(t_{n})\rangle$ $=\langle\phi,f^{\tau}(t_{n+1})-f^{\tau}(t_{n+1}^{-})\rangle+\langle\phi,f^{\tau}(t^{-}_{n+1})-f^{\tau}(t_{n})\rangle$
$=\tau\left\langle\phi,E_{\mathcal{F}}\left[f^{\tau}(t_{n+1}^{-})\right]\right\rangle+\int_{t_{n}}^{t_{n+1}}\left(\langle\nabla\phi\cdot\bm{b},f(s)\rangle+\frac{1}{2}\langle\nabla\cdot\bm{\Sigma}\nabla\phi,f(s)\rangle\right)ds$
$=\int_{t_{n}}^{t_{n+1}}\left(\langle\phi,E_{\mathcal{F}}[f^{\tau}(s)]\rangle+\langle\phi,T_{\mathcal{L}}[f^{\tau}(s)]\rangle\right)ds$
$\qquad+\int_{t_{n}}^{t_{n+1}}\langle\phi,E\left[f^{\tau}(t_{n+1}^{-})\right]-E\left[f^{\tau}(s)\right]\rangle ds\,.$
The first term corresponds exactly to the PDE operator, while the second is an error term coming from the fact that the evolutionary operator acts only at discrete times $t_{n}$ in the algorithm.
Next, thanks to the Lipschitz continuity of the evolutionary operator and of $f^{\tau}(s),s\in(t_{n},t_{n+1})$, we have
$\int_{t_{n}}^{t_{n+1}}|\langle\phi,E\left[f^{\tau}(t_{n+1}^{-})\right]-E\left[f^{\tau}(s)\right]\rangle|ds\leq C\int_{t_{n}}^{t_{n+1}}\|f^{\tau}(t_{n+1}^{-})-f^{\tau}(s)\|_{\mathrm{BL}}ds\leq C^{\prime}\tau^{2}\,.$
Altogether, one obtains the estimate
$\left|\langle\phi,f^{\tau}(t)-f^{\tau}(0)\rangle-\int_{0}^{t}\left(\langle\phi,E_{\mathcal{F}}[f^{\tau}(s)]\rangle+\langle\phi,T_{\mathcal{L}}[f^{\tau}(s)]\rangle\right)ds\right|$
$\leq\sum_{n=0}^{\lfloor t/\tau\rfloor+1}\int_{t_{n}}^{t_{n+1}\wedge t}\langle\phi,E\left[f^{\tau}(t_{n+1}^{-})\right]-E\left[f^{\tau}(s)\right]\rangle ds\leq C^{\prime}t\,\tau$
Then, by taking the limit $\tau\to 0$, one obtains for all $t\in[0,t_{\max}]$
$\langle\phi,f(t)-f(0)\rangle=\int_{0}^{t}\left(\langle\phi,E_{\mathcal{F}}[f(s)]\rangle+\langle\phi,T_{\mathcal{L}}[f(s)]\rangle\right)ds\,$,
and so $f\in C([0,t_{\max}],\mathcal{P}_{q}(\mathbb{R}^{d}))$ is a weak measure solution to the Cauchy problem (30).
Appendix C Proof of Theorem 4.1: convergence to the fittest solution
C.1 Proof of Theorem 10
We divide the proof in different steps. First, thanks to a bound on the solution’s energy, we show that the mass around the maximizer $h^{\star}$ does not vanish over the computational interval $0,t_{\max}]$. Then, we apply the quantitative Laplace principle (22) and the mean evolution (19) to show that it converges towards the fittest point $h^{\star}$.
Evolution of second moments.
We denote the energy, or second moment, of $\rho$ as $\textup{En}(\rho)=M_{2}(\rho)=\int|h|^{2}\rho(dh)$. Thanks to Assumption 9 and (21), we have
$M_{2}\left(G_{\alpha\overline{\mathcal{F}}}[\rho]\right)\leq b_{1}+b_{2}M_{2}(\rho)$
where $b_{1},b_{2}>0$ can be made independent of the selection pressure, by simply assuming, for instance, $\alpha>1$. Since the evolutionary operator consists of selection and convolution with a Gaussian kernel, we also have $M_{2}(E_{\alpha\overline{\mathcal{F}}}(\rho))\leq C(1+M_{2}(\rho))$ with $C=C(\sigma,b_{1},b_{2})>0$ a constant independent on $\alpha$. Similarly to Lemma 15, one can obtain an estimates over the the entire time horizon
$M_{2}(\rho(t))\leq C(1+M_{2}(\rho_{0}))e^{C^{\prime}t}\qquad t\in[0,t_{\max}]$
for a solution $\rho\in C([0,T],\mathcal{P}_{2}(\mathbb{R}^{d_{h}}))$ to the averaged PDE (13). We omit the details for brevity. This allows us to infer that there exists $\overline{M}>0$ such that
$\sup_{t\in[0,t_{\max}]}M_{2}(G_{\alpha\overline{\mathcal{F}}}[\rho(t)])\leq\overline{M}\qquad\textup{for all}\quad\alpha>1\,.$ (35)
Lemma 16 (Mass around maximizer)
Let $r>0$ be fixed and $\overline{\mathcal{F}}$ satisfy Assumption 9. For any given measure $\rho\in\mathcal{P}_{2}(\mathbb{R}^{d_{h}})$ with $M_{2}(G_{\alpha\overline{\mathcal{F}}}[\rho])\leq\overline{M}$, there exists $\delta=\delta(r,\overline{M})>0$ such that for
$\mu:=E_{\alpha\overline{\mathcal{F}}}[\rho]+\rho\qquad\textup{it holds}\qquad\mu(B(h^{\star},r))\geq\delta\,.$
Proof Let $R>\sqrt{\overline{M}}$, and such that $|h^{\star}|<R$. By Markov’s inequality, there exists $\delta_{R}\in(0,1)$ such that
$G_{\alpha\overline{\mathcal{F}}}[\rho](B(0,R))\geq 1-\frac{M_{2}\left(G_{\alpha\overline{\mathcal{F}}}[\rho]\right)}{R^{2}}\geq 1-\delta_{R}\,.$
Next, recall that after selection we perform mutation by convolution with a Gaussian measure: $\mu=K_{\sigma}*G_{\alpha\overline{\mathcal{F}}}[\rho]$ . Since the Gaussian kernel is not degenerate, there exists a density lower bound $\underline{K}_{R}>0$ such that for all $h\in B(0,R)$
$\int\chi_{B(h^{\star},r)}(h+\sigma\xi)K_{\sigma}(d\xi)\geq|B(h^{\star},r)|\underline{K}_{R}\,$,
where $\chi_{A}(\cdot)$ is the indicator function for a set $A$. It follows
$\mu(B(h^{\star},r))$ $=\iint\chi_{B(h^{\star},r)}(h+\sigma\xi)K_{\sigma}(d\xi)G_{\alpha\overline{\mathcal{F}}}[\rho](dh)$
$\geq\int_{\mathbb{R}^{d_{h}}}\int_{B(0,R)}\chi_{B(h^{\star},r)}(h+\sigma\xi)K_{\sigma}(d\xi)G_{\alpha\overline{\mathcal{F}}}[\rho](dh)$
$\geq|B(h^{\star},r)|\underline{K}_{R}G_{\alpha\overline{\mathcal{F}}}[\rho](B(0,R))$
$\geq|B(h^{\star},r)|\underline{K}_{R}\delta_{R}>0\,.$
Evolution of mass over $B(h^{\star},r)$.
Thanks to (35), we can apply Lemma 16 to the PDE solution $\rho(t)$, and obtain that for any $r>0$ there exists $\delta>0$ such that
$E_{\alpha\overline{\mathcal{F}}}[\rho(t)](B(h^{\star},r))$ $=\left(E_{\alpha\overline{\mathcal{F}}}[\rho(t)]+\rho(t)\right)(B(h^{\star},r))-\rho(t)(B(h^{\star},r))$
$\geq\delta-\rho(t,B(h^{\star},r))\,$,
leading to the estimate
$\frac{d}{dt}\rho(t,B(h^{\star},r))\geq\delta-\rho(t,B(h^{\star},r))\,.$
By Grönwall’s inequality, it holds
$\rho(t,B(h^{\star},r))\geq e^{-t}\rho_{0}(B(h^{\star},r))+(1-e^{-t})\delta\qquad\textup{for}\quad t\in[0,t_{\max}]\,.$
In particular, the mass over $B(h^{\star},r)$ is strictly positive during the entire interval $0,t_{\max}]$
Quantitative Laplace principle.
In the following, we denote the weighted mean as $\textup{m}^{\alpha}[\rho]:=\textup{m}\left(G_{\alpha\overline{\mathcal{F}}}[\rho(t)]\right)$ for simplicity. Let $tol>0$ be fixed. By the quantitative Laplace principle (22), for any $\rho\in\mathcal{P}_{2}(\mathbb{R}^{d_{h}})$, under the inverse continuity Assumption 9 (iii) we have that there exists $r_{tol}>0$ and $q_{tol}>0$ sufficiently small such that
$|\textup{m}^{\alpha}[\rho(t)]-h^{\star}|\leq\sqrt{\frac{{tol}}{8}}+\frac{\exp(-\alpha q_{tol})}{\rho(t,B(h^{\star},r_{tol}))}\int|h-h^{*}|\rho(dh)\,.$
Now, we apply the estimate for our solution $\rho(t)$, together with the previous estimates for the mass around the fittest solution, and energy bounds. In particular, we use that $M_{2}(\rho(t))\leq\overline{M}=C(1+M_{2}(\rho_{0}))e^{Ct_{\max}}$ for $t\in[0,t_{\max}]$, and Lemma 16 with $r_{tol}$ and $\overline{M}$. We get
$|\textup{m}^{\alpha}[\rho(t)]-h^{\star}|$ $\leq\sqrt{\frac{{tol}}{8}}+\frac{\exp(-\alpha q_{tol})}{\rho(t,B(h^{\star},r_{tol}))}\int|h-h^{*}|\rho(t,dh)$
$\leq\sqrt{\frac{{tol}}{8}}+\frac{\exp(-\alpha q_{tol})}{\rho_{0}(B(h^{\star},r_{tol}))\wedge\delta}2\left(|h^{\star}|+\sqrt{\overline{M}}\right)\,.$
Since $q_{tol}>0$, this estimate allows as to take $\alpha\gg 1$ sufficiently large such that the second term is of order $\sqrt{{tol}}$, and obtain
$|\textup{m}^{\alpha}[\rho(t)]-h^{\star}|\leq\sqrt{\frac{{tol}}{8}}+\sqrt{\frac{{tol}}{8}}=\sqrt{\frac{{tol}}{2}}\qquad\textup{for all}\quad t\in[0,t_{\max}]\,.$
By plugging this in the mean evolution, we get
$\frac{d}{dt}|\textup{m}[\rho(t)]-h^{\star}|^{2}$ $=2(\textup{m}[\rho(t)]-h^{\star})\cdot(\textup{m}^{\alpha}[\rho(t)]-\textup{m}[\rho(t)])$
$=-2|\textup{m}[\rho(t)]-h^{*}|^{2}+2(\textup{m}[\rho(t)]-h^{*})\cdot(\textup{m}^{\alpha}[\rho(t)]-\textup{m}[\rho(t)])$
$\leq-|\textup{m}[\rho(t)]-h^{*}|^{2}+|\textup{m}^{\alpha}[\rho(t)]-h^{\star}|^{2}$
$\leq-|\textup{m}[\rho(t)]-h^{*}|^{2}+\frac{{tol}}{2}\,.$
By Gröwnall’s inequality, we obtain the desired upper bound
$|\textup{m}[\rho(t)]-h^{\star}|^{2}\leq e^{-t}|\textup{m}[\rho_{0}]-h^{\star}|^{2}+(1-e^{-t})\frac{{tol}}{2}\,.$
By this estimate, one can directly see that if $t_{\max}$ is sufficiently large, as $t^{\star}$ in Theorem 10, then both terms in the right-hand-side can be made of order $tol/2$.
C.2 Additional proofs
Proof (Lemma 8) Fix $h\in H$. Under Assumption 7, the function
$\theta\longmapsto\mathcal{L}(\theta,h)-\mathcal{L}(\theta^{\star}(h),h)$
satisfies the assumptions of (Hasenpflug et al., 2024, Assumption 3.1) with $p=1$ and reference density $\pi_{0}\equiv 1$. The particular case where the solution is unique is discussed in (Hasenpflug et al., 2024, Remark 3.10). Recall $\mathbb{W}_{1}$ denotes the $L^{1}$-Wasserstein distance. By (Hasenpflug et al., 2024, Theorem 3.8)
$\mathbb{W}_{1}\bigl(\mu^{\infty}_{\mathcal{L}}(\cdot|h),\delta_{\theta^{\star}(h)}\bigr)\leq C\frac{1}{\sqrt{\beta}}.$
The constant can be be independent on $h$ thanks to the compactness of the hyperparameter space $H$ and the uniform estimate on the $\nabla^{2}\mathcal{L}(\theta^{\star}(h),h)$. Next, by Assumption 3, the fitness $\mathcal{F}$ is bounded and globally Lipschitz on $\mathbb{R}^{d_{\theta}}\times H$, and so is $e^{\mathcal{F}(\theta,h)}$. By Kantorovich–Rubinstein duality (Villani, 2009, Theorem 1.14), we have for some constant $C_{1}>0$ independent of $h$
$\left|\int e^{\mathcal{F}(\theta,h)}\,\mu^{\infty}_{\mathcal{L}}(d\theta|h)-e^{\mathcal{F}(\theta^{\star}(h),h)}\right|$ $=\left|\int e^{\mathcal{F}(\theta,h)}\,\mu^{\infty}_{\mathcal{L}}(d\theta|h)-\int e^{\mathcal{F}(\theta,h)}\,\delta_{\theta^{\star}(h)}(d\theta)\right|$
$\leq\|e^{\mathcal{F}(\cdot,h)}\|_{\textup{Lip}}\,\mathbb{W}_{1}\bigl(\mu^{\infty}_{\mathcal{L}}(\cdot|h),\delta_{\theta^{\star}(h)}\bigr)$
$\leq C_{1}\frac{1}{\sqrt{\beta}}.$
By Assumption 3, we have $\inf\mathcal{F}>0$ and so $\int e^{\mathcal{F}(\theta,h)}\,\mu^{\infty}_{\mathcal{L}}(d\theta|h)\geq e^{\inf\mathcal{F}}$ and $e^{\mathcal{F}(\theta^{\star}(h),h)}\geq e^{\inf\mathcal{F}}.$ Since the $\log$ function is globally Lipschitz on $e^{{\inf\mathcal{F}}},\infty)$, we obtain
$|\overline{\mathcal{F}}_{\beta}(h)-\mathcal{F}(\theta^{\star}(h),h)|$ $=\left|\log\int e^{\mathcal{F}(\theta,h)}\,\mu^{\infty}_{\mathcal{L}}(d\theta|h)-\log e^{\mathcal{F}(\theta^{\star}(h),h)}\right|$
$\leq e^{-{\inf\mathcal{F}}}\left|\int e^{\mathcal{F}(\theta,h)}\,\mu^{\infty}_{\mathcal{L}}(d\theta|h)-e^{\mathcal{F}(\theta^{\star}(h),h)}\right|$
$\leq C_{2}\frac{1}{\sqrt{\beta}},$
for some constant $C_{2}>0$ independent of $h$.
Appendix D Details of the Reinforcement Learning experiment
Q-learning.
In value-based reinforcement learning, an approach to maximise the expected cumulative reward is Q-learning (Watkins and Dayan, 1992). It aims to estimate the optimal action–value function $Q^{\ast}$, defined as
$Q^{\ast}(s,a)=\mathbb{E}\!\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\,\middle|\,s_{t}=s,\;a_{t}=a\right],$ (36)
where $\gamma\in(0,1)$ is a discount factor controlling the relative importance of future rewards.
The optimal action–value function satisfies the Bellman optimality equation
$Q^{\ast}(s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot\mid s,a)}\!\left[R(s,a)+\gamma\max_{a^{\prime}}Q^{\ast}(s^{\prime},a^{\prime})\right],$ (37)
which characterises $Q^{\ast}$ as the unique fixed point of the Bellman optimality operator.
Q-learning approximates this fixed point via an iterative, sample-based procedure, replacing the expectation in the Bellman equation with single-step transitions observed from the environment. The algorithm proceeds as follows:
  1. 1.
    At each time step, select an action according to a policy derived from the current estimate $Q$. A commonly used exploration strategy is the $p$-greedy policy: with probability $p$, a random action is selected (exploration), and with probability 1p1-p the greedy action
    $a(s)=\operatorname*{arg\,max}_{a^{\prime}}Q(s,a^{\prime})$ (38)
    is chosen (exploitation). The exploration rate $p$ is typically annealed from an initial value $p_{\mathrm{start}}$ to a final value $p_{\mathrm{end}}$, for example using an exponential schedule:
    $p_{\mathrm{next}}=p_{\mathrm{end}}+(p_{\mathrm{start}}-p_{\mathrm{end}})\exp\!\left(-\frac{\mathrm{total\ steps}}{p_{\mathrm{decay}}}\right).$ (39)
    We set $p_{\mathrm{start}}=1,p_{\mathrm{end}}=0.01$, and usually it starts with larger values of $p_{\mathrm{decay}}$ and follows an annealing schedule, so that in the beginning of the training it explores then exploits.
  2. 2.
    Execute the selected action and observe the reward $r$ and the next state $s^{\prime}$.
  3. 3.
    Update the action–value estimate according to
    $Q(s,a)\leftarrow Q(s,a)+\alpha\Bigl[r+\gamma\max_{a}Q(s^{\prime},a)-Q(s,a)\Bigr],$ (40)
    where $\alpha\in(0,1]$ is a learning rate, typically set to be decreased over time.
  4. 4.
    Set the current state to $s^{\prime}$ and repeat until a terminal condition is reached.
Under suitable conditions on exploration and learning rates, Q-learning converges to $Q^{\ast}$ for finite Markov decision processes.
Deep Q-Networks.
For problems with large or continuous state spaces, tabular Q-learning becomes infeasible due to the need to store $Q(s,a)$ for all state–action pairs. This motivates the use of function approximation, leading to the Deep Q-Network (DQN) algorithm (Mnih et al., 2015). In DQN, a neural network parameterised by $\theta$ is used to approximate the action–value function,
$Q(s,a;\theta)\approx Q^{\ast}(s,a).$
The network parameters are optimised by minimising the mean-squared temporal-difference error. At iteration $i$, the loss function is given by
$L_{i}(\theta_{i})=\mathbb{E}_{(s,a,r,s^{\prime})\sim U(D)}\left[\left(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta_{i}^{-})-Q(s,a;\theta_{i})\right)^{2}\right],$ (41)
where $\theta_{i}^{-}$ denotes the parameters of a target network, which is held fixed for several iterations to stabilise training and discount factor $\gamma$ was set to $0.99$ throughout. The expectation is approximated by sampling uniformly from a replay buffer $D=\{e_{1},\dots,e_{t}\}$, where each experience $e_{t}=(s_{t},a_{t},r_{t},s_{t+1})$ is stored during interaction with the environment. This mechanism, known as experience replay, reduces temporal correlations between samples and improves data efficiency, enabling stable training of deep value-based reinforcement learning algorithms.
We chose three hyperparameters, namely, learning rate of the Adam optimiser, the $p$-decay parameter $p_{\mathrm{decay}}$ from (39), and the experience replay batch size. Their respective ranges were set to $(10^{-5},10^{-2})$, $(500,5000)$, and $(32,128)$. During mutation, we scaled these domains to (1,1)(-1,1) before adding noise with the noise scale $\sigma=0.1$. This differs from the original PBT method and the Ray library, which typically perturb values by a multiplicative factor of $1.2$ or $0.8$.